### v1

# Teaching Human Behavior Improves Content Understanding Abilities Of LLMs

Abstract
Communication is defined as “Who says what to whom with what effect.” A message from a communicator generates downstream receiver effects, also known as behavior. Receiver behavior, being a downstream effect of the message, carries rich signals about it. Even after carrying signals about the message, the behavior data is often ignored while training large language models. We show that training LLMs on receiver behavior can actually help improve their content-understanding abilities. Specifically, we show that training LLMs to predict the receiver behavior of likes and comments improves the LLM’s performance on a wide variety of downstream content understanding tasks. We show this performance increase over 40 video and image understanding tasks over 23 benchmark datasets across both 0-shot and fine-tuning settings, outperforming many supervised baselines. Moreover, since receiver behavior, such as likes and comments, is collected by default on the internet and does not need any human annotations to be useful, the performance improvement we get after training on this data is essentially free-lunch. We release the receiver behavior cleaned comments and likes of 750k images and videos collected from multiple platforms along with our instruction-tuning data.
Equal Contribution. Contact ykumar@adobe.com for questions and suggestions.
Figure 1: The diagram depicts the five factors of communication in the context of an example YouTube video https://www.youtube.com/watch?v=p8ZYgb8Osek and where lies the free lunch. The receiver effect is not used while training Large Vision and Language Models. However, it contains many important signals that can help in understanding the content. The figure shows several comments containing temporal, cognitive, character, context, and user opinion information useful for understanding the video.
1 Introduction
Communication is defined by five factors: sender, message, channel, receiver, and behavior [71, 45, 46]. Lasswell encoded these five factors in the phrase, “Who says what to whom with what effect.” [71, 45, 46]. Human behavior occurs as a downstream artifact in the process of communication. Behavior is produced by the receiver as a response to the message sent by the sender. Being a downstream effect, behavior can help us infer important signals about the message itself. These signals, if properly harnessed, should be able to increase performance on the message understanding tasks popular in NLP and CV, like question answering, sentiment analysis, topic classification, etc. Despite this, behavior data is considered noise and is ignored while training large language models [15, 63] and also large vision and language models [54, 91]. In this paper, we explore this line of thought more.
Figure 2: Behavior-LLaVA is trained to answer behavioral questions like simulating user comments and likes on the video. The model, once trained, shows superior performance than LLaMA-Vid and other VLMs on content-related tasks like emotion recognition, action recognition, question answering, persuasion strategy classification, etc. The original video was showcased in SuperBowl-2024 and is posted on YouTube on the URL https://www.youtube.com/watch?v=OU7BJc96lI4. The video is titled “Perfect 10: The Kia big game commercial featuring the 2024 Kia EV9” by Kia America.
Humans produce two kinds of behavioral signals upon observing a message [11, 66]: perceptual signals and actions as behavior. Perceptual signals, like seeing, touching, and hearing, help a receiver primarily sense the world around her, ultimately guiding her actions. Actions are how a receiver acts on the outside world. The signals produced by the human receiver upon receiving a message carry information about the message itself (Fig. 1). For instance, the way a person’s eyes move while reading a book or how his heartbeat rises and falls upon watching a movie can help us extract useful information about the book or the movie. In this case, a rising heartbeat is often indicative of an exciting scene in the movie, and regressing while reading is indicative of important or confusing phrases [14]. Similarly, the actions a person performs after watching a movie, such as comments and likes, carry signals about the movie. Expanding on these ideas, prior literature has shown that harnessing perceptual signals, like eye movements, saliency, keystrokes, mouse movements, and FMRI, by modeling them together with content understanding tasks can improve both NLP and CV tasks. For instance, performance improvement is observed in tasks like visual and natural language question answering [62, 41, 74], text and image sentiment analysis [41, 10, 25], natural language inference [41], part-of-speech identification [8, 9], named entity recognition [31, 30], syntactic parsing [65], image captioning [21], visual object detection [77, 43].
While the initial studies show that perceptual signals have much promise for improving downstream content understanding, they have a few significant issues due to which integrating human perception has not seen wide adoption in training LLMs. The human processing signals can only be collected in lab settings requiring specialized lab equipment and are thus expensive to collect and are limited in number. For example, the largest datasets containing the human processing signals are SALICON [36] and Cheng et al. [19] for visual saliency (10k images each), CELER [12] containing 28k sentences and human eye movements of 365 participants reading them, Dundee corpus [38] containing eye movements over 20 news articles, and Dhakal et al. [22] containing keystroke patterns over 1.5k sentences. Clearly, these datasets, while making important contributions, do not scale to the level at which today’s large language models are trained (trillions of natural language and image tokens). Further, being restricted to lab settings also limits the types of demographics reachable by the lab, thus limiting their representativeness of human diversity. Lab settings also lead to people changing their behavior in response to their awareness of being observed, an effect commonly called the Hawthorne effect [59, 70, 58]. On the other hand, actions (the other type of behavioral signals produced by a human receiver) are collected at a large scale in the form of digital analytics. Examples of this kind of data are likes, views, shares, comments, and purchase histories on images, tweets, videos, webpages, and other kinds of media. Action data has a much broader representation than is possible in lab settings and is much cheaper to collect than using specialized lab equipment. At the same time, actions have not been much investigated in the literature for their potential to improve downstream content understanding. Further, they are much more noisier than perceptual signals collected in the lab, owing to their collection in real-world settings.
Therefore, in this paper, we make initial efforts to collect and understand digital analytics at scale with the aim of integrating them with LLMs to improve their downstream content understanding capabilities. We introduce methods for filtering and cleaning behavioral data and then propose tasks for large language and vision models, leading to improvements in language and visual content understanding tasks. For this, we look to Reddit and YouTube as two major sources of visual content and human behavior in the form of viewer comments, likes, and upvotes. From Reddit, we collect 5 million images and videos along with their upvotes and top-rated comments from two major subreddits (r/pics and r/videos). Similarly, from YouTube, we collect 2.2 million videos from 30000 channels along with their likes, views, and top user comments. After extensive filtering and cleaning, we are left with 730k samples of videos and images across the two platforms which we use for the next steps.
The Volkswagen ad titled begins with a group of three women seated excitedly in a Volkswagen Jetta, the driver sporting a wide smile as she grips the steering wheel. The voiceover sets a nostalgic tone, likening the Jetta to unforgettable first experiences like a first kiss or hearing indie rock for the first time—a symbol of newfound freedom and excitement. As the scenes unfold, the camera captures the trio cruising down winding roads, their laughter blending with the music and the wind tousling their hair. The atmosphere inside the car is one of camaraderie and adventure, with the Jetta serving as the backdrop to their shared moments of joy and spontaneity. Transitioning to a wider shot of the Jetta gliding along a scenic highway, surrounded by lush greenery, the ad evokes a sense of exploration and the open road. The visuals seamlessly blend modern-day cruising with vintage footage of classic Volkswagen vehicles, reflecting on the brand’s 75-year history in America, starting with the beloved Beetle. The ad concludes with the Volkswagen logo and the tagline ”An American Love Story”, encapsulating the enduring relationship between Volkswagen and its drivers across generations. This phrase serves as a tribute to Volkswagen’s 75-year history in America, beginning with the iconic Type 1 vehicles fondly known as ”The Beetle” Through its nostalgic narrative and captivating visuals, the teaser promises viewers an immersive journey into the essence of Volkswagen—a timeless icon that has been a part of countless cherished memories on the American road.
Figure 3: Dense caption generated by Behavior-LLaVA for the video of a Volkswagen ad. The original video is posted at URL: https://www.youtube.com/watch?v=kyuGXPNr-T0. The red-colored text highlights the most important aspects of the video captured by Behavior-LLaVA, demonstrating an understanding of aesthetics, characters, world knowledge, emotion, and spatial relationships. More such examples are given in Figs. 4, 5, 7, and Figs. 6, 8 for images and videos respectively.
In this captivating painting, a majestic tree takes center stage against a stunning backdrop. The sky above is a poetic split between night and day—a deep, starry blue on one side with a crescent moon, and a warm, radiant blend of yellows and pinks on the other, suggesting either a sunrise or sunset. The tree itself embodies the essence of transition, with one side adorned in lush green foliage and the other stripped bare, symbolizing the timeless dance of seasons and life’s perpetual transformations. This artwork not only captures the eye but also speaks to deeper truths. It invites us to contemplate the delicate balance between light and darkness, growth and dormancy, woven into the fabric of existence. The artist’s brushstrokes, rendered with a serene palette and gentle abstraction, evoke a sense of tranquility and introspection. This painting beautifully reminds us to embrace the fleeting beauty of each passing moment, whether bathed in daylight or cloaked in the quiet mystery of night.
Figure 4: Dense caption generated by Behavior-LLaVA for an artistic drawing of a tree.
After collecting user behavior over image and video content, next, we design tasks to teach large vision and language models (VLMs) to simulate user behavior. For this, we use an instruction fine-tuning format. Given a video or an image and the other metadata like time of post and channel, we ask the model to simulate user behavior of likes and comments. See Fig 2, Listing 1 for examples. We choose LLaMA-Vid [52] as our base model to teach it the user behavior. We call the resultant model Behavior-LLaVA (Large Language and Vision Assistant) [54]. Note that all the vision language models like LLaVA [54], MiniGPT-4 [91], Kosmos [32], BLIP [49], LLaMA-Vid [52] are trained on image-text instructions where text is some kind of annotations of the image, e.g., objects, object locations, captions, question-answers, etc. These instructions, since being sourced from either existing benchmark datasets such as MSCOCO [53] and Conceptual Captions [72] or being generated by stronger LLMs like GPT-4, require a significant amount of investment in terms of money for human annotation or generation through API. While building on top of their work, it is noteworthy that our model does not add any extra annotation or API cost to generate data. Rather, we rely on sampling data from digital analytics databases, which are anyway being collected for various reasons, and use them to improve VLM capabilities. Therefore, any potential improvements brought by integrating action data with VLMs do not require a significant capital investment.
The image features a pair of Nike Air Force 1 sneakers, which are a classic and iconic design. The sneakers are positioned in the center of the image, with the left shoe slightly in front of the right one, giving a sense of depth. They are predominantly white with green accents, including the Nike swoosh logo on the side and the outsole. The laces are white, and the midsole appears to be a light beige color. The background is a plain, light color, which contrasts with the sneakers and makes them stand out. Overlaid on the image is a stylized text that reads ”JUST DO IT” in a cursive, flowing script. The text is in a light green color that matches the accents on the sneakers, creating a cohesive color scheme. Below the main text, there is a smaller line of text that says ”New collection,” indicating that these sneakers are part of a new line or series from Nike. The overall style of the image is clean and modern, with a focus on the product. The use of the ”JUST DO IT” slogan, which is a well-known tagline for Nike, adds a layer of branding and recognition to the image. The text is designed to be eye-catching and to draw attention to the sneakers, which are the main subject of the image. The composition is balanced, with the sneakers centrally placed and the text evenly distributed around them.
Figure 5: Dense caption generated by Behavior-LLaVA for a Nike ad. The red-colored text highlights the most important aspects of the video captured by Behavior-LLaVA, demonstrating an understanding of aesthetics, characters, world knowledge, emotion, and spatial relationships.
We start with the trained LLaMA-Vid model. The LLaMA-Vid model uses two tokens to represent each video, which they call content and context tokens. While the context token encodes the overall image context based on user input, the content token encapsulates visual cues in each frame. For learning context tokens, the model uses attention queries that interact with previously generated image features in the designed attention module. To generate content tokens, the image features are average pooled. This dual-token strategy significantly reduces the number of tokens needed to represent videos, thus enabling the model to scale to longer (hour-long) videos. To better support hour-long videos, LLaMA-Vid was trained on a 9k movie-level conversation instruction set containing plot reasoning and detail understanding questions. Taking the base LLaMA-Vid model, we finetune it further on behavioral data. We teach the model a single instruction format given in Listing 1, where we teach it that given a media file (image or a video), it has to simulate the likes and comments on the media. We do this on a 730k instructions set created by us using Reddit and YouTube videos and images (§2). Next, we test the resultant model which we call Behavior-LLaVA, on a hierarchy of tasks in both zero-shot and finetuned settings. We include tasks based on low-level content understanding, like object and activity detection, and also on high-level understanding, like topic, sentiment, emotion, and persuasion classification. We also test the model on even higher-level tasks like memorability simulation and YouTube dislike/like simulation. We compare Behavior-LLaVA against the base model LLaMA-Vid and other state-of-the-art approaches reported on each test task (Tables 1-8).
We make the following contributions with this work:
1) We explore the idea of learning human behavior, resulting in better content understanding. We test this for action-level behavior data corresponding to receiver comments and likes. We collect a dataset consisting of 400k images and 330k videos along with their receiver comments and likes. Then, LLaMA-Vid is trained for the task of predicting receiver comments and likes given a media (a video or an image) (Listing 1). We show that using this simple task formulation and behavioral data collected in the wild, results in performance improvement over a hierarchy of tasks. We get improvements over the base LLaMA-Vid across 40 tasks over 23 benchmark datasets in both zero-shot and fine-tuned settings. We show this over low-level content understanding tasks like object and activity recognition and also over high-level tasks like topic and emotion detection. Further, interestingly, none of the tasks tested show a decrease in performance by training LLaMA-Vid on this additional task. Through this, we propose a scalable approach to increase the content understanding abilities of VLMs, requiring minimal cost and no architectural changes.
2) To disentangle the effect of training LLaVA-Vid on the additional 730k video and image samples from the effect of training on behavior data over those samples, we train LLaVA-Vid on 730k samples using the original task of LLaVA-Vid. We call this model Ad-LLaVA. We show that Ad-LLaVA shows slight performance improvements over LLaVA-Vid; however, Behavior-LLaVA still performs better than both Ad-LLaVA and LLaMA-Vid, thus highlighting the importance of behavior data and instruction fine-tuning on behavior data.
3) We also show an ablation of Behavior-LLaVA across different kinds of behavior. We try out the perception behavior of saliency prediction over images and two types of action-level behavior (comments and likes) over images and videos. We find that perception-level behavior does not result in significant performance improvements; however, action-level behavior shows improvements across all the tasks. We posit that one reason for this could be due to the scale for which action-level data is available (Table 9). While perception behavior is mostly collected in lab settings, action-level behavior data can be collected in a scalable manner automatically and cheaply.
2 Methodology
In this section, we introduce our approach to train Behavior-LLaVA. Since no publicly available corpus consists of behavior together with image and video content, we first introduce our instruction fine-tuning dataset, “Behavior-LLaVA Instruction Fine-Tuning dataset” (BLIFT). Next, we introduce our methodology to train Behavior-LLaVA. Finally, we report the results of testing Behavior-LLaVA’s capabilities on a hierarchy of tasks. The tasks cover low-level media understanding tasks like object and activity detection, high-level media understanding tasks like emotion, topic, and persuasion strategy classification.
2.1 BLIFT Dataset
Given the abundance of media and behavioral data and its accessibility, our data collection relies on two primary sources: Reddit and YouTube. These platforms share similarities in terms of hosting media content (images and videos) and user engagement metrics such as Reddit upvotes and comments, and YouTube likes, views, and comments. Here, we delineate the process involved in constructing the instruction fine-tuning dataset, which we term as the Behavior-LLaVA Instruction Fine-Tuning (BLIFT) dataset.
2.1.1 Data from Reddit
To collect a substantial corpus of diverse images and videos, we targeted two specific subreddits, namely r/pics and r/videos. Established over 15 years ago, these subreddits had a user base exceeding 20 million during the data collection period, with an average of over 5,000 online users concurrently. Notably, due to stringent content moderation guidelines [69, 2, 3, 4] and the exclusive focus of these subreddits on media content, they offer a rich variety of content devoid of thematic biases. Our data collection spans until January 2022, during which the activity on these subreddits witnessed a notable decline following several policy adjustments and user protests [29, 24].
To ensure data quality and relevance, we executed a series of filtering steps on the posts and comments from these subreddits. Initially, we excluded posts predating February 2015 from r/pics, coinciding with the implementation of a rule requiring images without digital/overlay text [2, 3]. This filtering step resulted in the exclusion of 3.1 million images and 2 million videos. Subsequently, considering the sustained popularity of both subreddits, with rankings within the top 20 since 2017 and consistent membership exceeding 20 million, we confined our dataset to posts from January 2018 onwards. This selection process yielded 1.4 million images and 1.1 million videos.
Further refinement of the dataset involved removing posts and comments marked as NSFW, BOT-generated, or [deleted], along with eliminating duplicate images and videos. This curation step reduced the dataset to 876,000 images and 983,000 videos. To address redundancy in comments, we excluded those comprising fewer than three words and employed TF-IDF-based deduplication with a similarity threshold of 0.6, determined through manual observations.
Following these steps, posts with fewer than two comments were filtered out, resulting in a dataset comprising 631,000 images and 397,000 videos. Additionally, videos exceeding a duration of 500 seconds were omitted, leaving 221,000 videos for analysis. Notably, images not directly hosted on Reddit were excluded due to scraping and copyright limitations. Similarly, for r/videos, only videos hosted on YouTube were considered. It is pertinent to mention that approximately 51% of YouTube videos collected during this period were either made private/unlisted or removed, resulting in 400,000 images and 80,000 videos, accompanied by 1.5 million and 312,000 comments, respectively. These comprehensive filtering steps ensured the construction of a diverse and relevant dataset for fine-tuning instruction-based models.
Listing 1: Behavior Instruction fine-tuning template for the video: https://www.youtube.com/watch?v=BKPQkjRF4yY
<SYSTEM>You are an AI visual assistant, you are given a detailed description of a media, followed by the actual media. Answer all questions as you are seeing the media<SYSTEM><USER>The video advertisement is titled Gatorade | Make Your Own Footsteps with Suni Lee for the brand Gatorade. The audio in the ad says ”[ASR HERE …]”.
Analyze this video deeply, then write scene by scene description of the video and answer the following questions. What percentage of viewers would like this video, and what would be the top-5 popular comments on this video?.
<video>.. </video>
<USER><ASSISTANT>
The scene by scene descriptions are:
Scene 1: The scene shows a woman looking off into the distance with an orange line going around the outside of the screen. The foreground colors of the scene are Black,Mud_Green,Gray,Dark_Gray, and the background colors are Dark_Brown,Black,Dark_Gray. The dominant tone of the scene is neutral. This scene is categorized by the tags cosmetic, eyebrow, face, girl, ponytail, stand, string, woman.
Scene 2: The scene shows a woman balancing on a skate board in a yard. The foreground colors of the scene are Black,Mud_Green,Dark_Gray,Olive, and the background colors are Black,Dark_Gray,Gray,Dark_Brown. The dominant tone of the scene is neutral. This scene is categorized by the tags athletic, balance, beam, car, girl, house exterior, hurdle, jog, legging, plank, rail, seesaw, woman, yard.
Scene 3: The scene shows a girl jumping over a wooden ramp in the backyard. The foreground colors of the scene are Black,Dark_Gray,Gray,Dark_Blue, and the background colors are Dark_Brown,Dark_Blue,Purple,Dark_Pink,Brown. The dominant tone of the scene is neutral. This scene is categorized by the tags arm, backyard, balance, diving board, girl, house exterior, ledge, plank, purple, rail, shirt, short, stand, stretch, woman, yard.
The video will be liked by 2.0%
1. Wow. Love it. Shes such an inspiration to the the next generation as well as everyone.
2. Inspiring and great story behind this commercial. Builds meaning and purpose in the hearts and mind of youth. Its been awhile since good meaningful ads have been made.
3. Shes an inspiration to the world thanks to her my niece learning Gymnastics hopefully someday, she is an inspiration to other youngers as how suni is an inspiration to everyone
4. Chills watching this. Such an inspiration
5. Yooooo, this is straight up !
<ASSISTANT>
2.1.2 Data from YouTube
Our data collection from YouTube begins with querying Wikidata [76] for YouTube IDs to compile a list of channels. Wikidata, derived from Wikipedia, provides a curated selection of renowned channels, automatically filtering out noisy videos commonly found in datasets collected from diverse sources like user-generated videos. This initial step yielded a dataset of 2.2 million videos spanning the period from 2018 to 2023, sourced from approximately 6,000 channels collected from Wikidata.
To refine the dataset, manual filtering was employed to exclude certain categories deemed less relevant for our purposes. These categories included music and songs, gaming content, non-English videos, sports commentary, anime, memes, channels with disabled comments sections, and news-related content. Furthermore, videos exceeding a duration of 60 seconds were omitted, and only videos with a substantial viewership, defined as greater than 10,000 views, were retained. We observed that these videos usually have less noisy comments and likes.
Subsequently, the top comments from each video, as ordered by YouTube (i.e., the most liked comments), were selected for inclusion in the dataset. To address redundancy in comments, a TF-IDF filter was applied with a threshold of 0.7, which proved effective in removing duplicate comments prevalent in YouTube data.
Comments were further filtered to include only those with a minimum of four words and a maximum of 100 words, ensuring a balance between relevance and conciseness. Additionally, to mitigate the presence of NSFW content, a vocabulary specific to NSFW terms [1] was employed to filter out inappropriate posts. On average, we finally get 3.1 comments per video, providing a substantial corpus of user-generated content for analysis. After applying these filtering steps, the dataset was reduced to 250,000 videos, ensuring a curated and relevant collection for subsequent analysis and model training.
2.2 Instruction Fine-Tuning LLaMA-Vid
After collecting Reddit and YouTube media and user behavior, we formulate instruction fine-tuning tasks for training LLaMA-Vid. In the training instruction, given the media content and automatic speech recognition if available, we ask the model to simulate the scene-by-scene description and user likes/views and top-5 comments. This instruction training template is given in Listing 1. To generate the instruction data, first, frames are sampled using the 30-degree rule [73, 5, 26], then the scene-by-scene description are obtained by concatenating automatically generated captions and tags from LLaVA-13B [54], colors and tone through Qin et al. [67]. This instruction format keeps the instructions similar to the instruction format for other VLMs like LLaVA [54], MiniGPT-4 [91], BLIP [48], LLaMA-Vid [52], etc., while additionally teaching the model to learn behavior. We keep the instruction fine-tuning template similar for both YouTube and Reddit. The complete instruction is given in Listing 1.
To disentangle the effect of training on additional data samples from the effect of training on behavioral data, we train LLaMA-Vid on BLIFT with the video and image verbalization and do not include receiver behavior of comments and likes. Then, the overall instruction template consists of scene-by-scene automatically generated verbalization similar to Listing 1 without the likes and comment simulation. We call the LLaMA-Vid fine-tuned on this data, Ad-LLaVA. We compare Behavior-LLaVA with Ad-LLaVA and LLaMA-Vid along with other state-of-the-art literature benchmarks on various tasks (Tables 1-8).
Listing 2: Perceptual Signal Instruction fine-tuning template for the image: http://farm6.staticflickr.com/5106/5670500150_e035dd2d30_z.jpg
<SYSTEM>You are an AI visual assistant. Answer all questions as you are seeing the media<SYSTEM><USER>The objects in this image in no particular order are car, dog, frisbee. Give me the order of saliency of these objects, start with the most salient object and end with the least salient object, each in a separate line. Give me the objects only and nothing else.
<image>
<ASSISTANT>
dog
frisbee
car
<ASSISTANT>
As an ablation experiment, we also try teaching the Behavior-LLaVA perceptual signals. For this, we take the largest perception signal dataset in the literature - Salicon10k [36]. It consists of 10,000 MS COCO images [18] with free-viewing eye gaze data collected through a novel mouse-based interface. The dataset has been widely used in many studies. We formulate two tasks using this data, (1) Salicon [Object]: estimating the saliency over the objects in the image and (2) Salicon [Region]: estimating the saliency over a region, where the regions tiles formed by breaking the image into a 3x3 grid. For both tasks we try to model two objectives, ranking and predicting, we found ranking to be much more effective. The instruction is given in Listings 2 and 3.
Listing 3: Perceptual Signal Instruction fine-tuning template for the image: http://farm6.staticflickr.com/5106/5670500150_e035dd2d30_z.jpg
<SYSTEM>You are an AI visual assistant. Answer all questions as you are seeing the media<SYSTEM><USER>Assume the given image is broken into a 3X3 grid the regions or tiles being named upper-left upper-center”, upper-right”, middle-left”, middle-center”, middle-right”, bottom-left”, bottom-center”, bottom-right”. Rank these regions or tiles based on their saliency, give me the line separated ranking of all regions in decreasing order.
<ASSISTANT>
middle-right
bottom-center
bottom-right
upper-center
upper-right
middle-center
upper-left
middle-left
bottom-left
We take the pretrained checkpoints of the LLaMA-VID 13B [51] model after the modality-alignment stage. We combine our instruction pairs from BLIFT with the original instruction tuning dataset consisting of 40K text conversations from ShareGPT, 625K single or multi-turn visual QA pairs, and 98K video QA pairs, all the modules except the Visual Encoder are kept frozen. We train the LLaMA-Vid checkpoints with their original sft mix along with BLIFT. We ablate different sampling ratios and found 1:1 to be empirically performing the best. We train the model for 2.2 epochs, keeping track of the 0-shot evaluation metrics and perplexity on comments in the eval set. For the best checkpoint, the perplexity on comments reduces from 6.22 to 3.05, and the $R^{2}$ on likes/views goes from -5.1 to 0.45.
3 Results and Discussion
In the experimental results, we aim to showcase the diverse and emergent capabilities of our Behavior-LLaVA model through quantitative numbers on various tasks and qualitative examples. These abilities include generating detailed image and video descriptions, emotion and sentiment analysis, question answering, video understanding tasks like scene and action detection. Additionally, we present the ability of Behavior-LLaVA to transfer learn on other behaviors like memorability of a video - both short-term and long-term.
Model scene way_speaking relationship like_ratio view_count director genre writer year Video4096-GPT-3.5 generated story + Flan-t5-xxl [13] 60.2 39.07 64.1 0.061 12.84 69.9 58.1 52.4 75.6 Video4096-GPT-3.5 generated story + GPT-3.5 classifier [13] 54.54 32.95 68.42 0.031 12.69 75.26 50.84 32.16 75.96 LLaMA-Vid + GPT-3.5 Generated Story 58.12 35.5 60.6 0.314 10.34 65.34 49.77 34.23 72.12 Ad-LLaVA 59.05 37.07 61.2 0.319 10.37 66.84 55.13 35.33 77.34 Behavior LLaVA + GPT-3.5 Generated Story 66.43 41.03 64.21 0.17 5.12 71.12 63.45 39.4 79.3 Improvement of Behavior LLaVA over LLaMA-Vid 10.48% 15.58% 9.62% 45.86% 50.48% 8.85% 27.49% 15.1% 9.96%
Table 1: Comparison of various models on the Long Video Understanding benchmark [79] consisting of 9 VQA tasks. We see that Behavior-LLaVA improves on LLaMA-Vid on 9/9 tasks with an average improvement of 21.49%. Further, it outperforms the state-of-the-art in 5/9 tasks.
Model Scene Object Action Event Attribute Concept Overall 0-shot LLaMA-VID 47.25 65.26 78.12 28.03 42.33 50.03 51.83 Ad-LLaVA 49.10 65.35 77.45 31.45 43.33 50.70 52.91 Behavior-LLaVA 52.03 65.33 77.95 32.66 45.67 51.20 54.14 Improvement of Behavior-LLaVA over LLaMA-Vid 10.12% 0.11% -0.22% 16.52% 7.89% 2.34% 4.46% 0-shot w/ story Video4096 - GPT-3.5 generated story + Flan-t5-xxl classifier [13] 59.66 98.89 98.96 38.42 67.76 86.99 75.12 Video4096 - GPT-3.5 generated story + GPT-3.5 classifier [13] 60.2 99.16 98.72 40.79 67.17 88.6 75.77 LLaMA-VID + Generated Story 60.3 99.92 99.01 39.33 66.66 87.33 75.425 Behavior-LLaVA + GPT3.5 generated story 60.4 99.89 98.23 40.97 67.23 88.33 75.84 Improvement of Behavior-LLaVA over LLaMA-Vid 0.16% -0.03% -0.78% 4.17% 0.85% 1.14% 0.55%
Table 2: Comparison of various models on the Holistic Video Understanding benchmark [23] consisting of 7 VQA tasks. We see that Behavior-LLaVA improves on LLaMA-Vid on 6/7 tasks with an average improvement of 5.88%. Further, it performs equivalent to the state-of-the-art in 6/7 tasks. State of the art is achieved by generating a story and asking Behavior-LLaVA to answer questions based on the generated story.
3.1 Evaluation
To test the effectiveness of Behavior-LLaVA, we conduct experiments involving 40 distinct tasks across 23 benchmark datasets. The diversity of tasks and datasets allows us to evaluate the performance and capabilities of Behavior-LLaVA thoroughly. Each of them is covered briefly next:
Method LLM MSVD-QA MSRVTT-QA ActivityNet-QA Acc Score Acc Score Acc Score FrozenBiLM [84] DeBERTa-V2 32.2 16.8 24.7 VideoLLaMA [87] Vicuna-7B 51.6 2.5 29.6 1.8 12.4 1.1 LLaMA-Adapter [88] LLaMA-7B 54.9 3.1 43.8 2.7 34.2 2.7 VideoChat [50] Vicuna-7B 56.3 2.8 45.0 2.5 26.5 2.2 Video-ChatGPT [56] Vicuna-7B 64.9 3.3 49.3 2.8 35.2 2.7 BT-Adapter [55] Vicuna-7B 67.5 3.7 57.0 3.2 45.7 3.2 LLaMA-VID Vicuna-7B 69.7 3.7 57.7 3.2 47.4 3.3 LLaMA-VID Vicuna-13B 70.0 3.7 58.9 3.3 47.5 3.3 Ad-LLaVA Vicuna-13B 70.0 3.7 59.0 3.3 47.4 3.3 Behaviour-LLaVA Vicuna-13B 70.1 3.7 59.2 3.4 47.5 3.3 Improvement of Behavior-LLaVA over LLaMA-Vid 0.14% 0% 0.5% 3% 0% 0%
Table 3: Comparison of various models on three conventional video question answering benchmarks consisting of question answers related to action understanding. The main goal of comparing on this benchmark is to show that Behavior-LLaVA does not perform worse on low-level understanding tasks like action recognition. We see that Behavior-LLaVA marginally improves on LLaMA-Vid on 2/3 benchmarks. Further, it performs equivalent to the state-of-the-art in 3/3 benchmarks.
  1. 1.
    Visual Question Answering (VQA): We evaluate the performance of visual question answering on the following benchmark datasets:
    • The Long-Video Understanding (LVU) benchmark by Wu and Krahenbuhl [79] comprises nine distinct tasks aimed at assessing long video comprehension, incorporating over 1000 hours of video content. These tasks encompass diverse aspects such as content understanding (including ‘relationship’, ‘speaking style’, ‘scene/place’), prediction of user engagement (‘YouTube like ratio’, ‘YouTube popularity’), and forecasting movie metadata (‘director’, ‘genre’, ‘writer’, ‘movie release year’). Evaluation of LVU tasks by Wu and Krahenbuhl [79] utilizes top-1 classification accuracy for content understanding and metadata prediction tasks, while Mean Squared Error (MSE) is employed for user engagement prediction tasks.
    • The Holistic Video Understanding (HVU) dataset, as introduced by Diba et al. [23], stands as the largest dataset for long video comprehension, comprising 572,000 samples. Encompassing a broad spectrum of semantic elements within videos, HVU tasks involve the classification of scenes, objects, actions, events, attributes, and concepts. Performance evaluation on HVU tasks, akin to the original work, is conducted using the mean average precision (mAP) metric on the validation set.
    • MSVD-QA and MSRVTT-QA: These datasets are based on Microsoft Research Video Description [17] and MSR-VTT corpora [82] and are extensively used in many video captioning and question-answering experiments. The MSVD-QA dataset has a total number of 1,970 video clips and 50,505 question-answer pairs. The MSRVTT-QA dataset contains 10K video clips and 243k question-answer pairs.
    • ActivityNet-QA [16] is a benchmark primarily for human activity understanding containing 849 hours of video, including 28,000 action instances.
  2. 2.
    Video and Image Understanding Benchmarks: We use a wide variety of tasks to evaluate video and image understanding: topic, emotion, and persuasion strategy classification, action and reason retrieval and generation, and emotions. We briefly introduce the benchmarks:
    • The advertisements dataset by Hussain et al. [33] contains 3,477 video advertisements and the corresponding annotations for emotion and topic tags and action-reason statements for each video. There are a total of 38 topics and 30 unique emotion tags per video. Further, we have 5 action-reason statements for each video for the action-reason generation task.
    • Persuasion strategy dataset [13] is a dataset consisting of 1002 video advertisements from popular brands and their persuasion strategy labels like social identity, anchoring and comparison, reciprocity, foot-in-the-door, etc.
      Training Model Topic Sentiment Persuasion Action Reason Clubbed All labels Random Random 2.63 3.37 14.3 8.37 3.34 3.33 Zero-shot VideoChat [50] 9.07 3.09 5.1 10.28 - - Video4096 - GPT-3.5 Generated Story + GPT-3.5 Classifier [13] 51.6 11.68 79.69 35.02 66.27 59.59 Video4096 - GPT-3.5 Generated Story + Flan-t5-xxl Classifier [13] 60.5 10.8 79.10 33.41 79.22 81.72 Video4096 - GPT-3.5 Generated Story + Vicuna Classifier [13] 22.92 10.8 67.35 29.6 21.39 20.89 Video4096 - Vicuna Generated Story + GPT-3.5 Classifier [13] 46.7 5.9 80.33 27.54 61.88 55.44 Video4096 - Vicuna Generated Story + Flan-t5-xxl Classifier [13] 57.38 9.8 76.60 30.11 77.38 80.66 Video4096 - Vicuna Generated Story + Vicuna Classifier [13] 11.75 10.5 68.13 26.59 20.72 21.00 LCBM [39] 42.17 7.08 58.83 32.83 39.55 27.91 LLaMA-VID w/ only video 10.11 3.42 5.75 12.32 29.61 24.11 LLaMA-VID w/ video + GPT-3.5 Story 42.72 11.05 64.02 32.07 37.76 42.33 Behavior-LLaVA w/ only video 22.65 11.13 60.04 13.39 42.66 33.33 Behavior-LLaVA w/ video + verbalization 46.34 11.7 64.13 33.33 52.06 52.03 Ad-LLaVA w/ video + GPT-3.5 story 51.16 11.33 68.03 33.11 43.26 51.45 Behavior-LLaVA w/ video + GPT-3.5 story 60.09 12.84 79.94 36.12 67.10 79.18 Improvement of Behavior-LLaVA over LLaMA-Vid 40.66% 16.2% 24.86% 12.62% 77.7% 87.05% Finetuned VideoMAE [75] 24.72 29.72 85.55 11.17 - - Hussain et al. [33] 35.1 32.8 - - 48.45 - Intern-Video [78] 57.47 36.08 86.59 5.47 6.8 - Video4096- Generated Story + Roberta Classifier [13] 71.3 33.02 84.20 64.67 42.96 39.09 LLaMA-VID w/ video + verbalization 59.13 32.11 79.15 50.93 50.32 30.13 LLaMA-VID w/ video + GPT-3.5 Story 63.11 35.01 84.15 55.01 57.11 45.73 Behavior-LLaVA w/ only video 58.03 22.72 84.41 26.23 59.33 51.45 Behavior-LLaVA w/ video + verbalization 68.32 33.92 85.93 64.72 70.89 75.34 Ad-LLaVA w/ video + GPT-3.5 story 66.34 36.24 84.09 58.31 68.15 78.15 Behavior-LLaVA w/ video + GPT-3.5 story 71.2 39.55 86.17 65.03 80.44 81.67 Improvement of Behavior-LLaVA over LLaMA-Vid 12.82% 12.97% 2.4% 18.21% 40.85% 78.59%
      Table 4: Comparison of various models on two video understanding benchmarks [33, 44] consisting of 5 tasks related to video advertisements understanding. The main goal of comparing on these benchmarks is to demonstrate Behavior-LLaVA’s understanding of complex videos. We see that Behavior-LLaVA improves on LLaMA-Vid on 5/5 tasks with an average improvement score of 43.18% in zero-shot and 27.64% in fine-tuned settings. Further, it outperforms the current state-of-the-art on 3/5 tasks in zero-shot and 5/5 in fine-tuned settings.
    • VideoEmotion-8 [7] dataset comprises 1,101 user-generated videos sourced from YouTube and Flickr, each containing a minimum of 100 videos per emotional category, as per Plutchik Wheel’s emotion model.
    • Ekman-6 [81] dataset is compiled from social websites, with each of its 1,637 videos labeled with a single emotion category based on Ekman’s psychological research.
    • CAER [47] dataset, sourced from TV shows, consists of 13,201 clips with an average sequence length of 90, each manually labeled with six basic emotions, aligning with the Ekman-6 dataset.
    • IAPSa [60] is a subset of IAPS, following the Mikels model with eight emotion categories such as amusement, awe, contentment, excitement, anger, disgust, fear, and sadness. It consists of 395 affective images, marking the first visual emotion dataset with discrete categories.
    • Emotion6 [64] features 1,980 images sourced from Flickr, each labeled by 15 annotators according to the Ekman model, covering six emotion categories: happiness, anger, disgust, fear, sadness, and surprise.
    • EmoSet [86] encompasses a total of 3.3 million images, including 118,102 from social networks and artistic sources, evenly distributed across various emotion categories. Based on the Mikels model, EmoSet is categorized into eight emotion categories.
    • Abstract [57] exclusively consists of color and texture combinations without recognizable objects. Differing from the IAPS dataset where emotions often stem from identifiable objects, the abstract paintings dataset was peer-rated via a web survey, with each image rated approximately 14 times. It comprises 228 images spanning eight categories similar to those in IAPS.
  3. 3.
    Image Dense Captioning: Literature image captioning datasets such as MS-COCO [18] reduce the inherently rich information and fine-grained semantics to simplistic captions, with very brief statements often focussing only on salient objects. This leads to a significant reduction in information content. Behavior data such as user comments help a model learn much more information such as object and material properties, world knowledge, emotion, character understanding, spatial relationships, aesthetics, etc. (see Fig. 1), enhancing the model’s captioning capability. Therefore, we design a captioning task to test this capability and compare it with respect to LLaMA-Vid and LLaVA-34B (a 2.5x larger model). Since we do not have ground truth for this task, following the LLM-as-a-judge paradigm, we use GPT-4V as the judge for all the models. GPT-4V is asked to evaluate the dense captions on three metrics: Correctness (Listing 4) evaluating the factuality and model hallucinations, Detail (Listing 5) evaluating the number and depth of details captured by the generated captions, and Quality (Listing 6) measuring the subjective quality of the concepts chosen to be highlighted by the captioning model and the arrangement, coherence, and the linking of various concepts.
  4. 4.
    Image and Video Memorability Simulation: Behavior-LLaVA is trained on behavior along with the media. To check if training on behavior helps in solving other behavior tasks [39], we test it over image and video memorability simulation. For this, we select seven benchmark datasets covering long-term and short-term memorability over images and videos.
    • Memento10k[61] is a short-term video memorability dataset comprising 10,000 video clips, with 900,000 human memory annotations recorded at various delay intervals. The video clips were, on average, 3s long.
    • VideoMem [20] is comprised of 10,000 soundless videos, each lasting 7 seconds, accompanied by memorability scores. Memorability was measured twice: first, shortly after viewing and again 24-72 hours later to capture both short-term and long-term memorability effects.
    • LaMem [40] dataset is a short-term image memorability dataset comprising of 60000 images. The dataset contains scene-centric images, object-centric images and other types such as images of art, images evoking certain emotions, and other user-generated images.
    • SUN [35] dataset is a short-term image memorability dataset comprising $2$ images that were sourced from the SUN database.
      Training Dataset Video Emotion-8 CAER Ekman-6 Random Random 12.5 14.28 16.67 0-Shot LLaMA-Vid 29.7 27.2 37.33 Behavior-LLaVA 41.35 51.0 49.33 Ad-LLaVA 29.8 27.3 37.66 Improvement of Behavior-LLaVA over LLaMA-Vid 39.22% 84.19% 32.14% Finetuned Zhao et al. [90] 54.5 78.3 55.3 Zhang et al. [89] 57.3 80.1 58.2 eMOTIONS [80] - - 53.12 Arevalo et al. [6] 53.7 77.3 54.2 Qiu et al. [68] 53.3 - 57.3 Xu et al. [81] 52.6 77.9 55.6 LLaMA-Vid 53.8 75.6 57.9 Ad-LLaVA 54.1 76.1 57.8 Behavior-LLaVA 56.9 79.3 58.4 Improvement of Behavior-LLaVA over LLaMA-Vid 5.76% 3.57% 0.86%
      Table 5: Comparison of various models on three video emotion understanding benchmarks (Video Emotion8 [37], CAER [47], Ekman-6 [81]). The main goal of comparing on these benchmarks is to demonstrate Behavior-LLaVA’s understanding of complex tasks like video emotions of long-form videos. We see that Behavior-LLaVA improves on LLaMA-Vid on 3/3 benchmarks with an average improvement score of 51.85% in zero-shot and 3.39% in fine-tuned settings. Further, it outperforms the current state-of-the-art on 3/3 benchmarks in zero-shot and 1/3 in fine-tuned settings.
    • MemCat [27] dataset is a short-term image memorability dataset comprising 10000 images. It consists of five broader memorability-relevant semantic categories (animal, sports, food, landscapes, vehicles), with 2K exemplars each, further divided into different subcategories (e.g., bear, pigeon, cat, etc. for animal). The images were sourced from existing image sets: ImageNet, COCO, Open Images Dataset, and SUN.
    • MediaEval [42] utilized publicly available links to short-form video clips, each averaging 6 seconds in duration, with both short-term and long-term memorability scores. Short-term memorability evaluations were conducted on videos viewed within the preceding few minutes, while long-term memorability assessments were based on videos viewed within the previous 24 to 72 hours.
    • LAMBDA [34] is a long-term memorability consisting of 2205 video ads collected over 1749 participants covering 276 brands. The average video length is 33 seconds and the videos are highly complex, consisting of audio, logos, fast-moving scenes, emotions, etc.
For Tables 1, 2, and 4, we follow the evaluation protocol of Video-4096 [13], for Table 3 we follow the evaluation protocol of LLaVA and LLaMA-VID. For Tables 5, 6, and 8, for 0-shot evaluation results, we use the logits of the next token from the given task vocabulary. For Table 8, we use the evaluation protocol of SI et al. [73].
In a heart-pounding and visually stunning trailer for Red Dead Redemption, we are thrust into the gritty world of the American frontier. The trailer opens with a voiceover, a chilling warning delivered with calm certainty: ”Listen to me, we don’t want to kill any of you… But trust me, we will.” Scenes flash by in quick succession, each more intense than the last. We see a lone figure, silhouetted against a setting sun, riding a magnificent horse through a sprawling, golden field. The rugged beauty of the landscape contrasts sharply with the impending sense of danger. Cut to a dimly lit saloon where a group of hardened men sit around a table, cards in hand, tension thick in the air. The voiceover continues, ”This whole thing is pretty much done. We’re more ghosts than people.” A flurry of action unfolds: a quick draw in a darkened room, bullets slicing through the air with deadly precision. The voiceover reminisces, ”Good old Dutch, my best friend… You know how we met? A pair of hucksters trying to rob each other… Back in ’78 or thereabouts.” The visuals intensify as we witness a robbery in progress, chaos erupting as masked figures burst into a bank. ”Ladies and gentlemen, this is a robbery,” declares one of the outlaws, setting the stage for a clash between lawlessness and order. Amidst the chaos, snippets of camaraderie emerge: ”Sons of Dutch. Makes us brothers.” But looming shadows of betrayal and regret cast doubt on these bonds. ”Sometimes, brothers make mistakes,” acknowledges the voiceover, acknowledging the complexities of loyalty and survival in this unforgiving world. The trailer crescendos with a crescendo of gunfights, horseback chases, and impassioned speeches. ”You’ll never change. I know that,” declares a voice, capturing the immutable nature of the human spirit amidst adversity. Throughout, a thematic motif resonates: ”You have to love yourself a fire.” The elemental symbolism underscores the primal nature of existence in a land where survival demands courage and cunning. As the trailer draws to a close, we’re left with the haunting refrain: ”Stay strong. Stay with me.” The screen fades to black, leaving us hungry for more of this epic tale set against the backdrop of the untamed West. With breathtaking visuals and evocative storytelling, this Red Dead Redemption trailer promises an unforgettable journey through a world where danger lurks around every corner, and the line between hero and outlaw blurs in the dust and shadows of the frontier.
Figure 6: Dense caption generated by Behavior-LLaVA for the video of the official trailer of the game Red Dead Redemption 2. The original video is posted at URL: https://www.youtube.com/watch?v=eaW0tYpxyp0.
The composition of the image presents a compelling narrative of a soldier immersed in a war-torn landscape. Positioned amidst a backdrop of explosive chaos and dense foliage, the soldier, adorned in traditional military gear, gazes directly at the viewer with a resolute demeanor. The soldier’s face is camouflaged, blending seamlessly with the olive-green helmet—a hallmark of battlefield attire. Surrounding this central figure, a tableau of activity unfolds: a tank looms in the background, its barrel skyward, suggesting recent action. Other soldiers, alert and vigilant, navigate the dense jungle terrain, underscoring the high stakes of the conflict. The foliage, lush yet foreboding, heightens the palpable tension of the scene. Foregrounded by the silhouette of another soldier’s helmeted head, the viewer is drawn into the heart of the action, evoking a sense of shared experience amid the perils of warfare. The realism of the depiction accentuates the emotional weight of the moment, capturing the essence of human resilience amidst the ravages of battle. Overall, the image transcends mere representation, offering a poignant reflection on the individual’s role in the broader narrative of war—imbued with suspense, authenticity, and a profound exploration of the human condition within conflict.
Figure 7: Dense caption generated by Behavior-LLaVA for a painting of a soldier. The model captures many qualitative aspects that are usually missed in common captioning tasks.
The scene is set at a packed stadium buzzing with excitement, the air thick with anticipation as Argentina and France face off in what promises to be an epic World Cup final. Lionel Messi steps up confidently to take a crucial penalty early in the match, the tension palpable as he eyes the goal. The commentator’s voice echoes through the stadium, ”He’s got the ground, he’s got a penalty! A heart beats… And Messi!” The crowd holds its breath as Messi strikes the ball cleanly, sending it into the back of the net. The stadium erupts with cheers as Argentina takes the lead, setting the stage for a thrilling encounter. As the game unfolds, both teams showcase their attacking prowess. Argentina’s Angel Di Maria emerges as a standout, capitalizing on a swift counterattack to score a spectacular goal. ”Alvarez, here’s McAllister, Di Maria is the spare man, and how Di Maria! Glorious goal! Argentine gold!” The Argentine fans go wild, waving their flags and chanting passionately. However, France fights back fiercely, earning a penalty of their own which they duly convert. The drama intensifies in the second half as France’s Kylian Mbappé dazzles with his speed and skill, scoring a breathtaking equalizer. ”Mbappé, to Lamp, Mbappé! Oh wow!” The match swings back and forth, with both teams pushing for victory. In the dying moments of extra time, the score is deadlocked at 3-3 when Argentina’s goalkeeper makes a crucial save, denying France a late winner. The match heads to a nail-biting penalty shootout. With nerves of steel, Argentina’s players step up and convert their penalties flawlessly, culminating in a decisive save by the goalkeeper. ”Argentina, champions of the world!” The stadium erupts once more, as Messi lifts the World Cup trophy amidst a sea of jubilant fans, marking a historic and unforgettable victory for Argentina
Figure 8: Dense caption generated by Behavior-LLaVA for the video of Argentina vs France FIFA World Cup Qatar 2022 Highlights. The original video is posted at URL: https://www.youtube.com/watch?v=zhEWqfP6V_w.
Training Models IAPSa-8 Abstract Emotion6 Emoset Random Random 12.5 12.5 16.67 12.5 0-shot GPT4-V 83.33 71.12 65.47 79.16 LLaMA-Vid 43.41 43.24 40.37 45.23 Ad-LLaVA 43.22 43.01 43.21 44.38 Behavior-LLaVA 57.97 64.21 49.71 50.38 Improvement of Behavior-LLaVA over LLaMA-Vid 33.54% 48.5% 23.14% 11.39% Finetuned MIDAN [83] 85.96 78.34 61.66 75.75 Stimuli-aware [85] - - 61.62 78.40 LLaMA-VID finetuned 84.93 71.23 62.87 80.31 Ad-LLaVA finetuned 85.13 71.16 62.66 79.88 Behavior-LLaVA finetuned 87.36 81.41 72.31 83.21 Improvement of Behavior-LLaVA over LLaMA-Vid 2.86% 14.29% 15.02% 3.61%
Table 6: Comparison of various models on four image emotion understanding benchmarks (IAPSa-8 [60] Abstract [57], Emotion6 [64], Emoset [86]). The main goal of comparing on these benchmarks is to demonstrate Behavior-LLaVA’s understanding of complex tasks like image emotions. We see that Behavior-LLaVA improves on LLaMA-Vid on 4/4 benchmarks with an average improvement score of 29.14% in zero-shot and 8.95% in fine-tuned settings. Further, it outperforms the current state-of-the-art on 4/4 benchmarks in the fine-tuned settings.
Model Correctness Detail Quality Average GPT4-V 8.4 8.5 8.4 8.43 LLaVA-1.6 (34B) 8.1 8.2 7.4 7.9 LLaMA-Vid (13B) 7.4 7.6 7.2 7.4 Ad-LLaVA (13B) 7.5 7.8 7.3 7.53 Behavior-LLaVA (13B) 7.3 8.1 7.9 7.76 Improvement of Behavior-LLaVA over LLaMA-Vid -1.3% 6.57% 9.72% 4.8%
Table 7: Comparison of various models on the image dense captioning task. The main goal of this task is to demonstrate Behavior-LLaVA’s image captioning ability. Despite not being explicitly trained on this task, Behavior-LLaVA performs better than both Ad-LLaVA and LLaMA-Vid on Detail and Quality aspects while losing marginally on correctness. On the aspects of detail and quality, it even outperforms the much larger model of LLaVA-1.6 (34B).
Training Models Image Datasets Video Datasets Lamem Memcat SUN Memento10k VideoMem MediaEval LAMBDA Human Consistency 0.68 0.78 0.75 0.73 0.61 - 0.61 Finetuned 10-shot in-context learning GPT-3.5 0.29 0.18 0.15 0.07 0.06 0.06 0.06 ViTMem [28] 0.71 0.65 0.63 0.56 0.51 - 0.08 Henry trained with 25% data [34] 0.56 0.64 0.59 0.62 0.49 0.32 0.28 Henry trained with 50% data [34] 0.65 0.68 0.67 0.69 0.55 0.44 0.40 Henry trained with 75% data [34] 0.71 0.75 0.73 0.74 0.62 0.49 0.47 Henry trained on all (combined) datasets [34] 0.72 0.79 0.76 0.72 0.60 0.48 0.52 Ad-LLaVA trained with 50% data 0.67 0.65 0.61 0.69 0.56 0.43 0.47 Behaviour LLaVA trained with 25% data 0.67 0.72 0.69 0.68 0.53 0.44 0.50 Behaviour LLaVA trained with 50% data 0.72 0.77 0.73 0.71 0.59 0.46 0.51 Behaviour LLaVA trained with 75% data 0.73 0.77 0.74 0.70 0.60 0.47 0.50 Behavior-LLaVA trained on all datasets 0.73 0.78 0.74 0.71 0.60 0.47 0.52 Improvement of Behavior-LLaVA over LLaMA-Vid (25% data) 19.64% 12.5% 16.95% 9.68% 8.16% 37.5% 78.57% Improvement of Behavior-LLaVA over LLaMA-Vid (50% data) 10.77% 13.26% 8.96% 2.90% 7.27% 4.54% 27.5% 0-shot LLaMA-Vid 0.13 0.11 0.05 0.03 0.05 0.02 0.05 Ad-LLaVA 0.14 0.13 0.06 0.06 0.07 0.04 0.13 Behavior-LLaVA 0.21 0.17 0.13 0.12 0.08 0.07 0.16 Improvement of Behavior-LLaVA over LLaMA-Vid 61.5% 54.5% 160% 300% 160% 350% 219%
Table 8: Comparison of various models on seven video and image memorability benchmarks (Memento10k [61], VideoMem [20], LaMem [40], SUN [35], MemCat [27], MediaEval [42], LAMBDA [34]). The main goal of comparing on these benchmarks is to demonstrate Behavior-LLaVA’s understanding of complex and high-level tasks like memorability simulation. We see that Behavior-LLaVA improves on LLaMA-Vid on 7/7 benchmarks with an average improvement score of 186.4% in zero-shot and 39% in fine-tuned settings after seeing 25% train data. Further, it performs similarly to the current state-of-the-art on 7/7 benchmarks in the fine-tuned settings while still seeing only 25% data.
Task MSRVTT-QA Emoset Comments likes/views Base 58.9 45.23 6.22 -0.1 Salicon [Region] 55.6 47.25 6.25 -0.07 Salicon [Object] 57.9 48.0 6.12 0.05 BLIFT[Likes/Views] 58.2 47.12 6.15 0.38 BLIFT[Comments] 59.0 49.58 3.02 0.19 BLIFT 59.2 50.38 3.05 0.40
Table 9: Ablation on using comments and/or perception signals from Salicon
3.2 Discussion
Tables 1, 2, and 3 contain the results for the visual question answering tasks, Tables 4, 5, 6 contain the results for video and image understanding tasks, Tables 7 contains the results for dense-captioning, and Table 8 contains the results for image and video memorability benchmarks. A common trend we observe across all the results is that Behavior-LLaVA performs better than the base model LLaMA-Vid and the finetuned model Ad-LLaVA on all tasks, especially in zero-shot settings. Interestingly, the performance gains remain even after fine-tuning on the task dataset (Tables 4, 5, 6).
The performance gains are smaller for the more low-level tasks of action and object recognition (Tables 1, 2, and 3), but much higher for the more high-level tasks of emotion understanding, sentiment analysis, persuasion strategy classification, and memorability simulation. This indicates that receiver behavior has richer signals for higher-level tasks. The gains are observed across both image and video benchmarks. We also observe that classification using a story generated by GPT-3.5 (following Bhattacharyya et al. [13]) results in better performance than only using the video (Tables 1, 2, 4).
Figures 4, 5, 7, and 3, 6, 8, show several randomly sampled qualitative examples for dense captions generated by Behavior-LLaVA over images and videos respectively. It can be noticed that despite not being explicitly trained for this task, the model performs quite well, picking up various artistic, cognitive, and object and material properties. From Table 7, while Behavior-LLaVA shows a decrease in correctness over LLaMA-Vid, it shows significant improvement in other aspects, including detail and quality. On these aspects, it even comes close to 2.5X larger models (LLaVA-1.6 (34B)).
Next, we compare the signals from behavioral data of perception and action. For this, we compare Behavior-LLaVA trained on the perception data of Salicon using two instruction formats with Behavior-LLaVA trained on BLIFT. Further, within BLIFT, we compare the learnings from the task of predicting likes/views with the task of predicting only comments. Table 9 contains the results. It can be noted from the table that training on just Salicon results in a performance decrease for the lower-level task of action recognition on MSRVTT-QA but improves performance on the higher-level task of Emotion recognition. However, the gains are smaller than those observed with training on BLIFT.
4 Conclusion
In this paper, we explore the idea of learning behavior leading to learning content better. Humans produce behavior in response to content. Hence, logically, behavior should contain signals about content, which, if used as a training task, should help in learning content better. We follow this line of thought and show that training large vision and language models on user behavior data of comments and likes collected from Reddit and YouTube leads to performance improvements across a wide variety of tasks. The gains are higher on higher-level tasks such as emotion recognition, persuasion strategy classification, and question answering and smaller on lower-level tasks like action and object recognition. Further, the gains remain even after fine-tuning the VLMs on those benchmarks, thus demonstrating the importance of learning behavior in understanding content better.
Appendix A GPT-4 Prompts
Listing 4: GPT-4V Prompt to calculate correctness of a image dense caption
You are a great critique for analyzing images and captions.
Assess the performance of a dense image captioning model based on the correctness of the captions generated.
Please assess the correctness of the provided caption in relation to the image. Consider whether the caption accurately identifies and describes the main subjects or objects depicted in the image. Assess whether the caption correctly interprets the relationships between elements within the image, such as actions, interactions, or spatial arrangements. Focus on the precision and accuracy of the information presented in the caption. Provide a score reflecting the level of correctness, ranging from 1 (low correctness) to 10 (high correctness).
Listing 5: GPT-4V Prompt to calculate detail of a image dense caption
You are a great critique for analyzing images and captions.
Assess the performance of a dense image captioning model based on the detail of the captions generated.
Evaluate the level of detail captured in the provided caption. Consider how well the caption describes specific attributes, features, or aspects of the image, including colors, shapes, textures, sizes, and any other relevant details. Assess whether the caption provides comprehensive information about the scene depicted in the image, covering both prominent and subtle elements. Pay attention to the depth and specificity of the details conveyed in the caption. Provide a score indicating the richness of detail, ranging from 1 (low detail) to 10 (high detail).
Listing 6: GPT-4V Prompt to calculate quality of a image dense caption
You are a great critique for analyzing images and captions.
Assess the performance of a dense image captioning model based on the quality of the captions generated.
Assess whether the caption is concise yet descriptive, providing meaningful and engaging information about the image. Evaluate the captions ability to evoke a clear mental image corresponding to the visual content. Additionally, consider if the caption is insightful or imaginative in its description. Provide a score reflecting the overall quality of the caption, ranging from 1 (low quality) to 10 (high quality).

### v3

@@
-Teaching Human Behavior Improves Content Understanding Abilities Of LLMs
+Veeky Baths [Uncaptioned image], Rajiv Ratn Shah [Uncaptioned image], Changyou Chen [Uncaptioned image], Balaji Krishnamurthy [Uncaptioned image]
 Abstract
@@
 Receiver behavior, being a downstream effect of the message, carries rich signals about it.
-Even after carrying signals about the message, the behavior data is often ignored while training large language models.
-We show that training LLMs on receiver behavior can actually help improve their content-understanding abilities.
-Specifically, we show that training LLMs to predict the receiver behavior of likes and comments improves the LLM’s performance on a wide variety of downstream content understanding tasks.
-We show this performance increase over 40 video and image understanding tasks over 23 benchmark datasets across both 0-shot and fine-tuning settings, outperforming many supervised baselines.
-Moreover, since receiver behavior, such as likes and comments, is collected by default on the internet and does not need any human annotations to be useful, the performance improvement we get after training on this data is essentially free-lunch.
-We release the receiver behavior cleaned comments and likes of 750k images and videos collected from multiple platforms along with our instruction-tuning data.
-Equal Contribution. Contact ykumar@adobe.com for questions and suggestions.
-Figure 1: The diagram depicts the five factors of communication in the context of an example YouTube video https://www.youtube.com/watch?v=p8ZYgb8Osek and where lies the free lunch.
+Even after carrying signals about the message, the behavior signal is often ignored while training vision language models.
+We show that training VLMs on receiver behavior can actually help improve their content-understanding abilities.
+We demonstrate that training VLMs to predict receiver behaviors, such as likes, comments, and replay graphs, which are available at scale, enhances the VLM’s performance across a broad range of downstream content understanding tasks.
+We show this performance increase over 6 types of behavior, 46 different tasks covering image, video, text and audio over 26 benchmark datasets across both 0-shot and fine-tuning settings, outperforming many supervised baselines on diverse tasks ranging from emotion recognition to captioning by upto 150%.
+We note that since receiver behavior, such as likes, comments, and replay graphs, is collected by default on the internet and does not need any human annotations to be useful, the performance improvement we get after training on this data is essentially free-lunch.
+We also release BLIFT, our Behaviour-LLaVA IFT dataset comprising of 730k images and videos with their receiver behavior collected from multiple platforms on which we train our models to achieve this.
+Figure 1: The diagram depicts the five factors of communication in the context of an example YouTube video https://www.youtube.com/watch?v=eT8hO4e2iTM and where lies the free lunch.
 The receiver effect is not used while training Large Vision and Language Models.
@@
 The figure shows several comments containing temporal, cognitive, character, context, and user opinion information useful for understanding the video.
+{NoHyper}Equal Contribution. Contact behavior-in-the-wild@googlegroups.com for questions and suggestions.
 1 Introduction
-Communication is defined by five factors: sender, message, channel, receiver, and behavior [71, 45, 46].
-Lasswell encoded these five factors in the phrase, “Who says what to whom with what effect.” [71, 45, 46].
-Human behavior occurs as a downstream artifact in the process of communication.
+Communication is defined by five factors: sender, message, channel, receiver, and behavior (Shannon & Weaver, 1949; Lasswell, 1948; 1971).
+Lasswell (1948) encoded these five factors in the phrase, “Who says what to whom with what effect.” Human behavior occurs as a downstream artifact in the process of communication.
 Behavior is produced by the receiver as a response to the message sent by the sender.
@@
 These signals, if properly harnessed, should be able to increase performance on the message understanding tasks popular in NLP and CV, like question answering, sentiment analysis, topic classification, etc.
-Despite this, behavior data is considered noise and is ignored while training large language models [15, 63] and also large vision and language models [54, 91].
+Despite this, behavior data is considered noise and is ignored while training large language models (Biderman et al., 2022; Penedo et al., 2023) and also large vision and language models (Liu et al., 2023a; Zhu et al., 2023).
 In this paper, we explore this line of thought more.
@@
 The video is titled “Perfect 10: The Kia big game commercial featuring the 2024 Kia EV9” by Kia America.
-Humans produce two kinds of behavioral signals upon observing a message [11, 66]: perceptual signals and actions as behavior.
+Humans produce two kinds of behavioral signals upon observing a message (Bertenthal, 1996; Prinz, 1997): perceptual signals and actions as behavior.
 Perceptual signals, like seeing, touching, and hearing, help a receiver primarily sense the world around her, ultimately guiding her actions.
@@
 1).
-For instance, the way a person’s eyes move while reading a book or how his heartbeat rises and falls upon watching a movie can help us extract useful information about the book or the movie.
-In this case, a rising heartbeat is often indicative of an exciting scene in the movie, and regressing while reading is indicative of important or confusing phrases [14].
-Similarly, the actions a person performs after watching a movie, such as comments and likes, carry signals about the movie.
+For instance, if a person’s heartbeat rises upon watching a movie scene, it can help us infer that perhaps the scene was an exciting scene (Dzedzickis et al., 2020).
+Similarly, regressing while reading is indicative of important or confusing phrases (Bicknell & Levy, 2011).
+In these cases, perception behavior helps us derive inferences about content.
+In a similar vein, the actions a person performs after watching a movie, such as comments and likes, carry signals about the movie (Fig.
+1, 2).
 Expanding on these ideas, prior literature has shown that harnessing perceptual signals, like eye movements, saliency, keystrokes, mouse movements, and FMRI, by modeling them together with content understanding tasks can improve both NLP and CV tasks.
-For instance, performance improvement is observed in tasks like visual and natural language question answering [62, 41, 74], text and image sentiment analysis [41, 10, 25], natural language inference [41], part-of-speech identification [8, 9], named entity recognition [31, 30], syntactic parsing [65], image captioning [21], visual object detection [77, 43].
+For instance, integration with perception signals causes performance improvement in tasks like visual and natural language question answering (Patro & Namboodiri, 2018; Khurana et al., 2023; Sood et al., 2020), text and image sentiment analysis (Khurana et al., 2023; Barrett et al., 2018; Fan et al., 2018), natural language inference (Khurana et al., 2023), part-of-speech identification (Barrett et al., 2016a; b), named entity recognition (Hollenstein et al., 2019; Hollenstein & Zhang, 2019), syntactic parsing (Plank, 2016), image captioning (Cornia et al., 2018), and visual object detection (Wang et al., 2018; Kruthiventi et al., 2016).
 While the initial studies show that perceptual signals have much promise for improving downstream content understanding, they have a few significant issues due to which integrating human perception has not seen wide adoption in training LLMs.
-The human processing signals can only be collected in lab settings requiring specialized lab equipment and are thus expensive to collect and are limited in number.
-For example, the largest datasets containing the human processing signals are SALICON [36] and Cheng et al.
-[19] for visual saliency (10k images each), CELER [12] containing 28k sentences and human eye movements of 365 participants reading them, Dundee corpus [38] containing eye movements over 20 news articles, and Dhakal et al.
-[22] containing keystroke patterns over 1.5k sentences.
+These perceptual signals can only be collected in lab settings requiring specialized lab equipment and are thus expensive to collect and thus are also limited in number.
+For example, the largest datasets containing the human processing signals are SALICON (Jiang et al., 2015) and Cheng et al.
+(2014) for visual saliency (10k images each), CELER (Berzak et al., 2022) and Dundee corpus (Kennedy et al., 2013) containing eye movements over 28k sentences and 20 news articles respectively, and Dhakal et al.
+(2018) containing keystroke patterns over 1.5k sentences.
 Clearly, these datasets, while making important contributions, do not scale to the level at which today’s large language models are trained (trillions of natural language and image tokens).
-Further, being restricted to lab settings also limits the types of demographics reachable by the lab, thus limiting their representativeness of human diversity.
-Lab settings also lead to people changing their behavior in response to their awareness of being observed, an effect commonly called the Hawthorne effect [59, 70, 58].
 On the other hand, actions (the other type of behavioral signals produced by a human receiver) are collected at a large scale in the form of digital analytics.
 Examples of this kind of data are likes, views, shares, comments, and purchase histories on images, tweets, videos, webpages, and other kinds of media.
-Action data has a much broader representation than is possible in lab settings and is much cheaper to collect than using specialized lab equipment.
+Action data has a much broader representation than is possible in lab settings, is available on more diverse content, and is much cheaper to collect than using specialized lab equipment.
 At the same time, actions have not been much investigated in the literature for their potential to improve downstream content understanding.
-Further, they are much more noisier than perceptual signals collected in the lab, owing to their collection in real-world settings.
-Therefore, in this paper, we make initial efforts to collect and understand digital analytics at scale with the aim of integrating them with LLMs to improve their downstream content understanding capabilities.
+Therefore, in this paper, we make initial efforts to collect and understand digital analytics at scale with the aim of integrating them with VLMs to improve their downstream content understanding capabilities.
 We introduce methods for filtering and cleaning behavioral data and then propose tasks for large language and vision models, leading to improvements in language and visual content understanding tasks.
-For this, we look to Reddit and YouTube as two major sources of visual content and human behavior in the form of viewer comments, likes, and upvotes.
-From Reddit, we collect 5 million images and videos along with their upvotes and top-rated comments from two major subreddits (r/pics and r/videos).
-Similarly, from YouTube, we collect 2.2 million videos from 30000 channels along with their likes, views, and top user comments.
+For this, we look to Reddit and YouTube as two major sources of visual content and human behavior in the form of viewer comments, likes, replay graphs, and upvotes.
+From Reddit, we collect 5 million images and videos along with their upvotes and top-upvoted comments from two major subreddits (r/pics and r/videos).
+Similarly, from YouTube, we collect 2.2 million videos from 30000 channels along with their likes, views, replay graphs, and top user comments.
 After extensive filtering and cleaning, we are left with 730k samples of videos and images across the two platforms which we use for the next steps.
+After collecting user behavior over image and video content, we design tasks to teach large vision and language models (VLMs) to simulate user behavior.
+For this, we use an instruction fine-tuning format.
+Given a video or an image and the other metadata like time of post and channel, we ask the model to simulate user behavior of likes and comments.
+See Fig 2, Listing 1 for examples.
+We choose LLaMA-Vid (Li et al., 2023b) as our base model to teach it the user behavior.
+We call the resultant model Behavior-LLaVA (Large Language and Vision Assistant) (Liu et al., 2023a).
+We test Behavior-LLaVA on a diverse variety of tasks, evaluating its capabilities on image, video, text, and audio understanding tasks.
+We compare Behavior-LLaVA against its base model, LLaMA-Vid, and other supervised baselines.
+Further, to show the impact of behavior, we train another version of LLaMA-Vid, where we train it on the same set of videos and images as Behavior-LLaVA but do not include behavior information.
+We call this model Ad-LLaVA.
+Figure 3: Behaviour-LLava achieves much higher zero-shot performance compared to Ad-LLaVA and the base model LLaMA-VID across a diverse suite of image, video, and audio benchmarks.
+We make the following contributions with this work:
+1) Behavior-LLaVA Instruction Fine-Tuning: We explore the idea of learning human behavior, resulting in better content understanding.
+We test this for action-level behavior data such as receiver comments, likes, and replay graphs.
+We collect a dataset called BLIFT, consisting of 400k images and 330k videos, along with their receiver behavior.
+Then, LLaMA-Vid is trained for the task of predicting receiver comments and upvotes given a media (a video or an image) (Listing 1).
+We show that using this simple task formulation over behavioral data collected in the wild, results in performance improvement over a hierarchy of tasks.
+We get improvements over the base LLaMA-Vid across 46 tasks over 26 benchmark datasets in both zero-shot and fine-tuned settings.
+We show this over low-level content understanding tasks like object and activity recognition and also over high-level tasks like topic and emotion detection.
+Through this, we propose a scalable approach to increase the content understanding abilities of VLMs, requiring minimal cost and no architectural changes.
+2) AdLLaVA: Disentangling the effect of content and behaviour: To disentangle the effect of training LLaMA-Vid on additional image and video data from the effect of training on behavior data, we train LLaMA-Vid on BLIFT’s videos and images without including behavior.
+We call this model Ad-LLaVA.
+We show that Ad-LLaVA shows equivalent performance as its base model LLaVA-Vid; however, Behavior-LLaVA performs better than both Ad-LLaVA and LLaMA-Vid, thus highlighting the importance of behavior data and instruction fine-tuning on behavior data.
+3) Perception vs Action:We also show an ablation of Behavior-LLaVA across different kinds of behavior.
+We try out the perception behavior of saliency prediction over images and five types of action-level behavior over images and videos.
+We find that perception-level behavior does not result in significant performance improvements; however, action-level behavior shows improvements across all the tasks.
+We posit that one reason for this could be due to the scale for which action-level data is available (Table 11).
+While perception behavior is mostly collected in lab settings, action-level behavior data is diverse, can be collected in a scalable manner automatically and cheaply.
+2 Methodology
+In this section, we introduce our approach to train Behavior-LLaVA.
+Since no publicly available corpus consists of behavior together with image and video content, we first introduce our instruction fine-tuning dataset, “Behavior-LLaVA Instruction Fine-Tuning dataset” (BLIFT).
+Next, we introduce our methodology to train Behavior-LLaVA.
+Finally, we report the results of testing Behavior-LLaVA’s capabilities on a hierarchy of tasks.
+The tasks cover low-level media understanding tasks like object and activity detection, high-level media understanding tasks like emotion, topic, and persuasion strategy classification.
+2.1 BLIFT Dataset
+Given the abundance of media and behavioral data and its accessibility, our data collection relies on two primary sources: Reddit and YouTube.
+These platforms share similarities in terms of hosting media content (images and videos) and providing user engagement metrics in the form of Reddit upvotes and comments, and YouTube likes, views, comments, and replay graphs.
+Here, we delineate the process involved in constructing the instruction fine-tuning dataset, which we term as the Behavior-LLaVA Instruction Fine-Tuning (BLIFT) dataset.
+2.1.1 Data from Reddit
+To collect a substantial corpus of diverse images and videos, we targeted at two specific subreddits, namely r/pics and r/videos.
+Established over 15 years ago, these subreddits had a user base exceeding 20 million during the data collection period, with an average of over 5,000 online users concurrently.
+Notably, due to stringent content moderation guidelines (Reddit, Inc., 2024; red, 2024a; b; c) and the exclusive focus of these subreddits on media content, they offer a rich variety of content devoid of thematic biases.
+Our data collection spans until January 2022, during which the activity on these subreddits witnessed a notable decline following several policy adjustments and user protests (Hern, 2023; Economist, 2023).
+To ensure data quality and relevance, we executed a series of filtering steps on the posts and comments from these subreddits.
+Initially, we excluded posts predating February 2015 from r/pics, coinciding with the implementation of a rule requiring images without digital/overlay text (red, 2024a; b).
+This filtering step resulted in the exclusion of 3.1 million images and 2 million videos.
+Subsequently, considering the sustained popularity of both subreddits, with rankings within the top 20 since 2017 and consistent membership exceeding 20 million, we confined our dataset to posts from January 2018 onwards.
+This selection process yielded 1.4 million images and 1.1 million videos.
+Further refinement of the dataset involved removing posts and comments marked as NSFW, BOT-generated, or [deleted], along with eliminating duplicate images and videos.
+This curation step reduced the dataset to 876,000 images and 983,000 videos.
+To address redundancy in comments, we excluded those comprising fewer than three words and employed TF-IDF-based deduplication with a similarity threshold of 0.6, determined through manual observations.
+Following these steps, posts with fewer than two comments were filtered out, resulting in a dataset comprising 631,000 images and 397,000 videos.
+Additionally, videos exceeding a duration of 500 seconds were omitted, leaving 221,000 videos for analysis.
+Notably, images not directly hosted on Reddit were excluded due to scraping and copyright limitations.
+Similarly, for r/videos, only videos hosted on YouTube were considered.
+It is pertinent to mention that approximately 51% of YouTube videos collected during this period were either made private/unlisted or removed, resulting in 400,000 images and 80,000 videos, accompanied by 1.5 million and 312,000 comments, respectively.
+These comprehensive filtering steps ensured the construction of a diverse and relevant dataset for fine-tuning instruction-based models.
+Listing 1: Behavior Instruction fine-tuning template for the video: https://www.youtube.com/watch?v=BKPQkjRF4yY
+<SYSTEM>You are an AI visual assistant. You are given a detailed description of a media, followed by the actual media. Answer all questions as if you are seeing the media.</SYSTEM>
+<USER>
+The video advertisement is titled "Gatorade | Make Your Own Footsteps with Suni Lee" for the brand Gatorade.
+The audio in the ad says "[ASR HERE ...]".
+Analyze this video deeply, then write scene by scene description of the video and answer the following questions.
+What percentage of viewers would like this video, and what would be the top-5 popular comments on this video?
+What would the replay graph values for each scene be?
+<video>...</video>
+</USER>
+<ASSISTANT>
+The scene-by-scene descriptions are:
+Scene 1: The scene shows a woman looking off into the distance with an orange line going around the outside of the screen.
+The foreground colors of the scene are black, mud green, gray, dark gray, and the background colors are dark brown, black, dark gray.
+The dominant tone of the scene is neutral.
+This scene is categorized by the tags: cosmetic, eyebrow, face, girl, ponytail, stand, string, woman.
+Scene 2: The scene shows a woman balancing on a skateboard in a yard.
+The foreground colors of the scene are black, mud green, dark gray, olive, and the background colors are black, dark gray, gray, dark brown.The dominant tone of the scene is neutral.
+This scene is categorized by the tags: athletic, balance, beam, car, girl, house exterior, hurdle, jog, legging, plank, rail, seesaw, woman, yard.
+Scene 3: The scene shows a girl jumping over a wooden ramp in the backyard.
+The foreground colors of the scene are black, dark gray, gray, dark blue, and the background colors are dark brown, dark blue, purple, dark pink, brown.
+$>$ BEHAVIOR $<$
+The video will be liked by 2.0%
+1. "Wow. Love it. Shes such an inspiration to the next generation as well as everyone."
+2. "Inspiring and great story behind this commercial. Builds meaning and purpose in the hearts and minds of youth. Its been a while since good, meaningful ads have been made."
+3. "Shes an inspiration to the world. Thanks to her,my niece is learning gymnastics. Hopefully someday, she is an inspiration to others as Suni is an to everyone"
+4. "Chills watching this. Such an inspiration."
+5. "Yooooo, this is straight up!"
+The replay values for each scene would be:
+Scene 1: 0.06
+Scene 2: 0.23
+Scene 3: 0.38
+</ASSISTANT>
+Figure 4: Behavior Instruction fine-tuning template for the video: https://www.youtube.com/watch?v=BKPQkjRF4yY
+2.1.2 Data from YouTube
+Our data collection from YouTube begins with querying Wikidata (Vrandečić & Krötzsch, 2014) for YouTube IDs to compile a list of channels.
+Wikidata, derived from Wikipedia, provides a curated selection of renowned channels, automatically filtering out noisy videos commonly found in datasets collected from diverse sources like user-generated videos.
+This initial step yielded a dataset of 2.2 million videos spanning the period from 2018 to 2023, sourced from approximately 6,000 channels collected from Wikidata.
+To refine the dataset, manual filtering was employed to exclude certain categories deemed less relevant for our purposes.
+These categories included music and songs, gaming content, non-English videos, sports commentary, anime, memes, channels with disabled comments sections, and news-related content.
+Furthermore, and only videos with a substantial viewership, defined as greater than 10,000 views, were retained.
+We observed that these videos usually have less noisy comments and likes.
+Subsequently, the top comments from each video, as ordered by YouTube (i.e., the most liked comments), were selected for inclusion in the dataset.
+To address redundancy in comments, a TF-IDF filter was applied with a threshold of 0.7, which proved effective in removing duplicate comments prevalent in YouTube data.
+Comments were further filtered to include only those with a minimum of four words and a maximum of 100 words, ensuring a balance between relevance and conciseness.
+Additionally, to mitigate the presence of NSFW content, a vocabulary specific to NSFW terms (ldn, ) was employed to filter out inappropriate posts.
+On average, we finally get 3.1 comments per video, providing a substantial corpus of user-generated content for analysis.
+After applying these filtering steps, the dataset was reduced to 250,000 videos, ensuring a curated and relevant collection for subsequent analysis and model training.
+2.2 Instruction Fine-Tuning LLaMA-Vid
+After collecting Reddit and YouTube media and user behavior, we formulate instruction fine-tuning tasks for training LLaMA-Vid.
+In the training instruction, given the media content and automatic speech recognition if available, we ask the model to simulate the scene-by-scene description and user likes/views and top-5 comments.
+This instruction training template is given in Listing 1.
+To generate the instruction data, first, frames are sampled using the 30-degree rule (SI et al., 2023; Arev et al., 2014; Friedman & Feldman, 2004), then the scene-by-scene description are obtained by concatenating automatically generated captions and tags from LLaVA-13B (Liu et al., 2023a), colors and tone through Qin et al.
+(2020).
+This instruction format keeps the instructions similar to the instruction format for other VLMs like LLaVA (Liu et al., 2023a), MiniGPT-4 (Zhu et al., 2023), BLIP (Li et al., 2022), LLaMA-Vid (Li et al., 2023b), etc., while additionally teaching the model to learn behavior.
+We keep the instruction fine-tuning template similar for both YouTube and Reddit.
+The complete instruction is given in Listing 1.
+We start with the trained LLaMA-Vid model.
+The LLaMA-Vid model uses two tokens to represent each frame in the video, which they call content and context tokens.
+While the context token encodes the overall image context based on user input, the content token encapsulates visual cues in each frame.
+For learning context tokens, the model uses attention queries that interact with previously generated image features in the designed attention module.
+To generate content tokens, the image features are average pooled.
+This dual-token strategy significantly reduces the number of tokens needed to represent videos, thus enabling the model to scale to longer (hour-long) videos.
+To better support hour-long videos, LLaMA-Vid was trained on a 9k movie-level conversation instruction set containing plot reasoning and detail understanding questions.
+Taking the base LLaMA-Vid model, we finetune it further on behavioral data.
+In our experiments we observe that the sampling ratio of BLIFT and IFT datasets is an important hyperparameter.
+We track 4 zero-shot metrics, likes/views, comments perplexity, and empirically find the best results with 1:1 ratio for 2.2 epochs (see Fig 11 and Table 12).
+For the best checkpoint, the perplexity on comments reduces from 6.22 to 3.05, and the R2 on likes/views goes from -5.1 to 0.45
+We combine 730k instruction pairs from BLIFT with the original instruction tuning dataset consisting of 40K text conversations from ShareGPT, 625K single or multi-turn visual QA pairs, and 98K video QA pairs; all the modules except the Visual Encoder are kept frozen.
+We ablate on multiple sampling ratios from BLIFT.
+We train the LLaMA-Vid checkpoints with their original SFT mix along with BLIFT.
+We ablate different sampling ratios and found 1:1 to be empirically performing the best.
+We train the model for 2.2 epochs, keeping track of the 0-shot evaluation metrics and perplexity on comments in the eval set Fig 11 and Table 12 show the ablations on different sampling ratios and epochs of training.
+For the best checkpoint, the perplexity on comments reduces from 6.22 to 3.05, and the $R^{2}$ on likes/views goes from -5.1 to 0.45.
+AdLLaVA to show the impact of behavior data: To disentangle the effect of training on additional data samples from the effect of training on behavioral data, we train LLaMA-Vid on BLIFT with the video and image verbalization and do not include receiver behavior.
+Then, the overall instruction template consists of scene-by-scene automatically generated verbalization similar to Listing 1 without the likes and comment simulation.
+We call the LLaMA-Vid fine-tuned on this data, Ad-LLaVA.
+We compare Behavior-LLaVA with Ad-LLaVA and LLaMA-Vid along with other state-of-the-art literature benchmarks on various tasks (Tables 1-4).
+Impact of filtering steps on performance: We use various filtering steps in our data pipeline, including NSFW filters, time filters, bot filters, and gaming and news video filters.
+To quantify the impact of our filtering steps on the final performance, we compare the performance from training on unfiltered data with filtered data (Figure 11).
+The figure shows the benefit of our data filtering process.
+While training on unfiltered BLIFT also improves the result over baseline performance of Llama-Vid but data filtering adds more improvement on top of it
+Ablation with perceptual behavior: As an ablation experiment, we also try teaching the Behavior-LLaVA perceptual signals.
+For this, we take the largest perception signal dataset in the literature - Salicon10k (Jiang et al., 2015).
+It consists of 10,000 MS COCO images (Chen et al., 2015) with free-viewing eye gaze data collected through a novel mouse-based interface.
+The dataset has been widely used in many studies.
+We formulate two tasks using this data, (1) Salicon [Object]: estimating the saliency over the objects in the image and (2) Salicon [Region]: estimating the saliency over a region, where the regions tiles formed by breaking the image into a 3x3 grid.
+For both tasks we try to model two objectives, ranking and predicting, we found ranking to be much more effective.
+The instruction are given in Listings 5 and 6.
+Model scene way_speaking relationship like_ratio view_count director genre writer year Video4096-GPT-3.5 generated story + Flan-t5-xxl 60.2 39.07 64.1 0.061 12.84 69.9 58.1 52.4 75.6 Video4096-GPT-3.5 generated story + GPT-3.5 classifier 54.54 32.95 68.42 0.031 12.69 75.26 50.84 32.16 75.96 LLaMA-Vid + GPT-3.5 Generated Story 58.12 35.5 60.6 0.314 10.34 65.34 49.77 34.23 72.12 Ad-LLaVA 59.05 37.07 61.2 0.319 10.37 66.84 55.13 35.33 77.34 Behavior LLaVA + GPT-3.5 Generated Story 66.43 41.03 64.21 0.17 5.12 71.12 63.45 39.4 79.3 Improvement of Behavior LLaVA over LLaMA-Vid 10.48% 15.58% 9.62% 45.86% 50.48% 8.85% 27.49% 15.1% 9.96%
+Table 1: Comparison of various models on the Long Video Understanding benchmark (Wu & Krahenbuhl, 2021) consisting of 9 VQA tasks.
+We see that Behavior-LLaVA improves on LLaMA-Vid on 9/9 tasks with an average improvement of 21.49%.
+Further, it outperforms the state-of-the-art in 5/9 tasks.
+3 Results and Discussion
+In the experimental results, we aim to showcase the diverse and emergent capabilities of our Behavior-LLaVA model through quantitative numbers on various tasks and qualitative examples.
+These abilities include generating detailed image and video descriptions, emotion and sentiment analysis, question answering, video understanding tasks like scene and action detection.
+Additionally, we present the ability of Behavior-LLaVA to transfer learn on other behaviors like memorability of a video - both short-term and long-term.
+3.1 Evaluation
+To test the effectiveness of Behavior-LLaVA, we conduct experiments involving 46 distinct tasks across 26 benchmark datasets.
+The diversity of tasks and datasets allows us to evaluate the performance and capabilities of Behavior-LLaVA thoroughly.
+Each of them is covered briefly next:
+  1. 1.
+    Visual Question Answering (VQA): We evaluate the performance of visual question answering on the following benchmark datasets:
+    • The Long-Video Understanding (LVU) benchmark by Wu & Krahenbuhl (2021) comprises nine distinct tasks aimed at assessing long video comprehension, incorporating over 1000 hours of video content.
+These tasks encompass diverse aspects such as content understanding (including relationship, speaking style, scene/place), prediction of user engagement (YouTube like ratio, YouTube popularity), and movie metadata (director, genre, writer, movie release year).
+    • The Holistic Video Understanding (HVU) dataset by Diba et al.
+(2020) stands as the largest dataset for long video comprehension, comprising 572,000 samples.
+Encompassing a broad spectrum of semantic elements within videos, HVU tasks involve the classification of scenes, objects, actions, events, attributes, and concepts.
+Performance evaluation on HVU tasks is conducted using the mean average precision (mAP) metric on the validation set.
+    • We also use MSVD-QA, MSRVTT-QA (Chen & Dolan, 2011; Xu et al., 2016b), and ActivityNet-QA (Caba Heilbron et al., 2015) datasets. Their description is given in Appendix A.
+  2. 2.
+    Video and Image Understanding Benchmarks: We use a wide variety of tasks to evaluate video and image understanding: topic, emotion, and persuasion strategy classification, action and reason retrieval and generation, and emotions.
+We briefly introduce the benchmarks:
+    • The advertisements dataset by Hussain et al.
+(2017) contains 3,477 video advertisements and the corresponding annotations for emotion and topic tags and action-reason statements for each video.
+There are a total of 38 topics and 30 unique emotion tags per video.
+Further, we have 5 action-reason statements for each video for the action-reason generation task.
+    • Persuasion strategy dataset (Bhattacharyya et al., 2023) is a dataset consisting of 1002 video advertisements from popular brands and their persuasion strategy labels like social identity, anchoring and comparison, reciprocity, foot-in-the-door, etc.
+      Training Model Topic Sentiment Persuasion Action Reason Clubbed All labels Random Random 2.63 3.37 14.3 8.37 3.34 3.33 Zero-shot VideoChat (Li et al., 2023a) 9.07 3.09 5.1 10.28 - - Video4096 - GPT-3.5 Generated Story + GPT-3.5 Classifier 51.6 11.68 79.69 35.02 66.27 59.59 LCBM (Khandelwal et al., 2024) 42.17 7.08 58.83 32.83 39.55 27.91 LLaMA-VID w/ only video 10.11 3.42 5.75 12.32 29.61 24.11 LLaMA-VID w/ video + GPT-3.5 Story 42.72 11.05 64.02 32.07 37.76 42.33 Behavior-LLaVA w/ only video 22.65 11.13 60.04 13.39 42.66 33.33 Behavior-LLaVA w/ video + verbalization 46.34 11.7 64.13 33.33 52.06 52.03 Ad-LLaVA w/ video + GPT-3.5 story 51.16 11.33 68.03 33.11 43.26 51.45 Behavior-LLaVA w/ video + GPT-3.5 story 60.09 12.84 79.94 36.12 67.10 79.18 Improvement of Behavior-LLaVA over LLaMA-Vid 40.66% 16.2% 24.86% 12.62% 77.7% 87.05% Finetuned Video4096- Generated Story + Roberta Classifier 71.3 33.02 84.20 64.67 42.96 39.09 LLaMA-VID w/ video + verbalization 59.13 32.11 79.15 50.93 50.32 30.13 LLaMA-VID w/ video + GPT-3.5 Story 63.11 35.01 84.15 55.01 57.11 45.73 Behavior-LLaVA w/ only video 58.03 22.72 84.41 26.23 59.33 51.45 Behavior-LLaVA w/ video + verbalization 68.32 33.92 85.93 64.72 70.89 75.34 Ad-LLaVA w/ video + GPT-3.5 story 66.34 36.24 84.09 58.31 68.15 78.15 Behavior-LLaVA w/ video + GPT-3.5 story 71.2 39.55 86.17 65.03 80.44 81.67 Improvement of Behavior-LLaVA over LLaMA-Vid 12.82% 12.97% 2.4% 18.21% 40.85% 78.59%
+      Table 2: Comparison of various models on two video understanding benchmarks (Hussain et al., 2017; Kumar et al., 2023) consisting of 5 tasks related to video advertisements understanding.
+The main goal of comparing on these benchmarks is to demonstrate Behavior-LLaVA’s understanding of complex videos.
+We see that Behavior-LLaVA improves on LLaMA-Vid on 5/5 tasks with an average improvement score of 43.18% in zero-shot and 27.64% in fine-tuned settings.
+Further, it outperforms the current state-of-the-art on 3/5 tasks in zero-shot and 5/5 in fine-tuned settings.
+Full results are presented in the Table 6.
+    • For emotion analysis, we use VideoEmotion-8 (Asur & Huberman, 2010), Ekman-6 (Xu et al., 2016a), CAER (Lee et al., 2019), IAPSa (Mikels et al., 2005), Emotion6 (Peng et al., 2015), EmoSet (Yang et al., 2023), and Abstract (Machajdik & Hanbury, 2010) datasets.
+A brief description for each of them is given in Appendix A.
+  3. 3.
+    Image Dense Captioning: Literature image captioning datasets such as MS-COCO (Chen et al., 2015) reduce the inherently rich information and fine-grained semantics to simplistic captions, with very brief statements focussing only on salient objects.
+Behavior data such as user comments help a model learn much more information such as object and material properties, world knowledge, emotion, character understanding, spatial relationships, aesthetics, etc.
+(see Fig.
+1), enhancing the model’s captioning capability.
+Therefore, we design a captioning task to test this capability and compare it with respect to LLaMA-Vid and LLaVA-34B (a 2.5x larger model).
+Since we do not have ground truth for this task, following the LLM-as-a-judge paradigm, we use GPT-4V as the judge for all the models.
+GPT-4V is asked to evaluate the dense captions on three metrics: Correctness (Listing 2) evaluating the factuality and model hallucinations, Detail (Listing 3) evaluating the number and depth of details captured by the generated captions, and Quality (Listing 4) measuring the subjective quality of the concepts chosen to be highlighted by the captioning model and the arrangement, coherence, and the linking of various concepts.
+  4. 4.
+    Image and Video Memorability Simulation: Behavior-LLaVA is trained on behavior along with the media.
+To check if training on behavior helps in solving other behavior tasks (Khandelwal et al., 2024), we test it over image and video memorability simulation.
+For this, we select seven benchmark datasets covering long-term and short-term memorability over images and videos: LaMem (Khosla et al., 2015), SUN (Isola et al., 2011), and MemCat (Goetschalckx & Wagemans, 2019) for images and Memento10k(Newman et al., 2020), VideoMem (Cohendet et al., 2019), MediaEval (Kiziltepe et al., 2021), and LAMBDA (I et al., 2024) for videos.
+We briefly cover each of them in Appendix A.
+  5. 5.
+    Modalities other than videos and images: Behavior-LLaVA, built on top of LLaMA-Vid and fine-tuned using BLIFT, is pretrained and fine-tuned on image and video datasets.
+To test if behavior data can improve the results on other modalities as well, we test Behavior-LLaVA’s performance on two tasks across audio and text modalities (Table 13).
+For audio, we evaluate on the audio summarization task (Han et al., 2023) and for text, we evaluate on the IMDB sentiment benchmark (Maas et al., 2011).
+    Training Dataset Video Emotion-8 CAER Ekman-6 Random Random 12.5 14.28 16.67 0-Shot LLaMA-Vid 29.7 27.2 37.33 Behavior-LLaVA 41.35 51.0 49.33 Ad-LLaVA 29.8 27.3 37.66 Improvement of Behavior-LLaVA over LLaMA-Vid 39.22% 84.19% 32.14% Finetuned Zhao et al.
+(2020) 54.5 78.3 55.3 Zhang et al.
+(2023c) 57.3 80.1 58.2 eMOTIONS (Wu et al., 2023) - - 53.12 Arevalo et al.
+(2017) 53.7 77.3 54.2 Qiu et al.
+(2020) 53.3 - 57.3 Xu et al.
+(2016a) 52.6 77.9 55.6 LLaMA-Vid 53.8 75.6 57.9 Ad-LLaVA 54.1 76.1 57.8 Behavior-LLaVA 56.9 79.3 58.4 Improvement of Behavior-LLaVA over LLaMA-Vid 5.76% 3.57% 0.86%
+    Table 3: Comparison of various models on three video emotion understanding benchmarks (Video Emotion8 (Jiang et al., 2014), CAER (Lee et al., 2019), Ekman-6 (Xu et al., 2016a)).
+The main goal of comparing on these benchmarks is to demonstrate Behavior-LLaVA’s understanding of complex tasks like video emotions of long-form videos.
+We see that Behavior-LLaVA improves on LLaMA-Vid on 3/3 benchmarks with an average improvement score of 51.85% in zero-shot and 3.39% in fine-tuned settings.
+Further, it outperforms the current state-of-the-art on 3/3 benchmarks in zero-shot and 1/3 in fine-tuned settings.
+For Tables 1, 10, and 2, we follow the evaluation protocol of Video-4096 (Bhattacharyya et al., 2023), for Table 7 we follow the evaluation protocol of LLaVA and LLaMA-VID.
+For Tables 3, 8, and 4, for 0-shot evaluation results, we use the logits of the next token from the given task vocabulary.
+For Table 4, we use the evaluation protocol of SI et al.
+(2023).
+Training Models Image Datasets Video Datasets Lamem Memcat SUN Memento10k VideoMem MediaEval LAMBDA Human Consistency 0.68 0.78 0.75 0.73 0.61 - 0.61 Finetuned 10-shot in-context learning GPT-3.5 0.29 0.18 0.15 0.07 0.06 0.06 0.06 ViTMem (Hagen & Espeseth, 2023) 0.71 0.65 0.63 0.56 0.51 - 0.08 Henry trained with 25% data (I et al., 2024) 0.56 0.64 0.59 0.62 0.49 0.32 0.28 Henry trained with 50% data (I et al., 2024) 0.65 0.68 0.67 0.69 0.55 0.44 0.40 Henry trained with 75% data (I et al., 2024) 0.71 0.75 0.73 0.74 0.62 0.49 0.47 Henry trained on all (combined) datasets (I et al., 2024) 0.72 0.79 0.76 0.72 0.60 0.48 0.52 Ad-LLaVA trained with 50% data 0.67 0.65 0.61 0.69 0.56 0.43 0.47 Behaviour LLaVA trained with 25% data 0.67 0.72 0.69 0.68 0.53 0.44 0.50 Behaviour LLaVA trained with 50% data 0.72 0.77 0.73 0.71 0.59 0.46 0.51 Behaviour LLaVA trained with 75% data 0.73 0.77 0.74 0.70 0.60 0.47 0.50 Behavior-LLaVA trained on all datasets 0.73 0.78 0.74 0.71 0.60 0.47 0.52 Improvement of Behavior-LLaVA over LLaMA-Vid (25% data) 19.64% 12.5% 16.95% 9.68% 8.16% 37.5% 78.57% Improvement of Behavior-LLaVA over LLaMA-Vid (50% data) 10.77% 13.26% 8.96% 2.90% 7.27% 4.54% 27.5% 0-shot LLaMA-Vid 0.13 0.11 0.05 0.03 0.05 0.02 0.05 Ad-LLaVA 0.14 0.13 0.06 0.06 0.07 0.04 0.13 Behavior-LLaVA 0.21 0.17 0.13 0.12 0.08 0.07 0.16 Improvement of Behavior-LLaVA over LLaMA-Vid 61.5% 54.5% 160% 300% 160% 350% 219%
+Table 4: Comparison of various models on seven video and image memorability benchmarks (Memento10k (Newman et al., 2020), VideoMem (Cohendet et al., 2019), LaMem (Khosla et al., 2015), SUN (Isola et al., 2011), MemCat (Goetschalckx & Wagemans, 2019), MediaEval (Kiziltepe et al., 2021), LAMBDA (I et al., 2024)).
+The main goal of comparing on these benchmarks is to demonstrate Behavior-LLaVA’s understanding of complex and high-level tasks like memorability simulation.
+We see that Behavior-LLaVA improves on LLaMA-Vid on 7/7 benchmarks with an average improvement score of 186.4% in zero-shot and 39% in fine-tuned settings after seeing 25% train data.
+Further, it performs similarly to the current state-of-the-art on 7/7 benchmarks in the fine-tuned settings while still seeing only 25% data.
+3.2 Discussion
+Tables 1, 10, and 7 contain the results for the visual question answering tasks, Tables 2, 3, 8 contain the results for video and image understanding tasks, Tables 9 contains the results for dense-captioning, Table 4 contains the results for image and video memorability benchmarks, and Table 13 contains the results for the audio and text tasks.
+A common trend we observe across all the results is that Behavior-LLaVA performs better than the base model LLaMA-Vid and the finetuned model Ad-LLaVA on all tasks, especially in zero-shot settings.
+In fact, Ad-LLaVA performs very similar to LLaMA-Vid itself.
+This shows that BLIFT adds meaningful signals on an average rather than noise to the model.
+Interestingly, the performance gains remain even after fine-tuning on the task dataset (Tables 2, 3, 8).
+The performance gains are relatively smaller for low-level tasks of action and object recognition (Tables 10, and 7), but much higher for the more high-level tasks of emotion understanding, sentiment analysis, persuasion strategy classification, and memorability simulation, longform-video understanding and other sub-tasks of Table 10.
+This indicates that receiver behavior has richer signals for higher-level tasks, infact fine-tuned Behaviour-LLaVA models outperform GPT4-V on image emotion recognition.
+The gains are observed across both image and video benchmarks.
+We also observe that classification using a story generated by GPT-3.5 (following Bhattacharyya et al.
+(2023)) results in better performance than only using the video (Tables 1, 10, 2).
+To evaluate the generalizability of behavior data across modalities, we extend our evaluation to audio summarization and text sentiment analysis and observe improvements of 19.5% (Table 13).
+Figures 6, 7, 10, and 5, 8, 9, show several randomly sampled qualitative examples for dense captions generated by Behavior-LLaVA over images and videos respectively.
+It can be noticed that despite not being explicitly trained for this task, the model performs quite well, picking up various artistic, cognitive, and object and material properties.
+From Table 9, while Behavior-LLaVA shows a decrease in correctness over LLaMA-Vid, it shows significant improvement in other aspects, including detail and quality.
+On these aspects, it even comes close to 2.5X larger models (LLaVA-1.6 (34B)).
+Next, in Table 11 we compare the signals from behavioral data of perception and action.
+For this, we compare Behavior-LLaVA trained on BLIFT and Behaviour-LLaVA trained on Salicon salient regions and objects.
+Further, within BLIFT, we compare the performance from predicting singled out behaviours including likes/views, titles, comments.
+It can be noted that training on just Salicon results in a performance decrease for the lower-level task of action recognition (MSRVTT-QA) but improves on the higher-level task of Emotion recognition.
+However, the gains are smaller than those observed with training on BLIFT.
+Model Summarization (3 Shot) Moment Retrieval Highlight Detection BLEU ROUGE METEOR mAP (avg) mAP Behavior-LLaVA[Replay-Graphs] 11.2 19.1 25.3 35.2 34.9 Behavior-LLaVA[Mem-Recalls] 11.9 20.3 26.9 34.9 33.1 Behavior-LLaVA 10.6 18.3 22.7 33.1 32.7 LLaMA-VID 9.1 16.5 20.3 30.3 32.4 Video-ChatGPT 5.0 14.0 19.7 - -
+Table 5: Improvement on downstream content understanding tasks by introducing more behaviour signals.
+Brackets [] denote the new behaviour that we include.
+Replay graphs (Khandelwal et al., 2024).
+Mem-Recalls (I et al., 2024) Evaluation done on Multi-shot video summarization (Han et al., 2023) and MomentDETR (Lei et al., 2021)
+4 Conclusion
+In this paper, we explore the idea of learning behavior leading to learning content better.
+Humans produce behavior in response to content.
+Hence, logically, behavior should contain signals about content, which, if used as a training task, should help in learning content better.
+We follow this line of thought and show that training large vision and language models on user behavior data of comments and likes collected from Reddit and YouTube leads to performance improvements across a wide variety of tasks.
+The gains are higher on higher-level tasks such as emotion recognition, persuasion strategy classification, and question answering and smaller on lower-level tasks like action and object recognition.
+Further, the gains remain even after fine-tuning the VLMs on those benchmarks, thus demonstrating the importance of learning behavior in understanding content better.
+Appendix
+Listing 2: GPT-4V Prompt to calculate correctness of a image dense caption
+You are a great critique for analyzing images and captions.
+Assess the performance of a dense image captioning model based on the correctness of the captions generated.
+Please assess the correctness of the provided caption in relation to the image.
+Consider whether the caption accurately identifies and describes the main subjects or objects depicted in the image.
+Assess whether the caption correctly interprets the relationships between elements within the image, such as actions, interactions, or spatial arrangements.
+Focus on the precision and accuracy of the information presented in the caption.
+Provide a score reflecting the level of correctness, ranging from 1 (low correctness) to 10 (high correctness).
+Listing 3: GPT-4V Prompt to calculate detail of a image dense caption
+You are a great critique for analyzing images and captions.
+Assess the performance of a dense image captioning model based on the detail of the captions generated.
+Evaluate the level of detail captured in the provided caption.
+Consider how well the caption describes specific attributes, features, or aspects of the image, including colors, shapes, textures, sizes, and any other relevant details.
+Assess whether the caption provides comprehensive information about the scene depicted in the image, covering both prominent and subtle elements.
+Pay attention to the depth and specificity of the details conveyed in the caption.
+Provide a score indicating the richness of detail, ranging from 1 (low detail) to 10 (high detail).
+Listing 4: GPT-4V Prompt to calculate quality of a image dense caption
+You are a great critique for analyzing images and captions.
+Assess the performance of a dense image captioning model based on the quality of the captions generated.
+Assess whether the caption is concise yet descriptive, providing meaningful and engaging information about the image.
+Evaluate the captions ability to evoke a clear mental image corresponding to the visual content.
+Additionally, consider if the caption is insightful or imaginative in its description.
+Provide a score reflecting the overall quality of the caption, ranging from 1 (low quality) to 10 (high quality).
+Listing 5: Perceptual Signal Instruction fine-tuning template for the image: http://farm6.staticflickr.com/5106/5670500150_e035dd2d30_z.jpg
+<SYSTEM>You are an AI visual assistant.
+Answer all questions as you are seeing the media<SYSTEM><USER>The objects in this image in no particular order are car, dog, frisbee.
+Give me the order of saliency of these objects, start with the most salient object and end with the least salient object, each in a separate line.
+Give me the objects only and nothing else.
+<image>
+<ASSISTANT>
+dog
+frisbee
+car
+<ASSISTANT>
+Listing 6: Perceptual Signal Instruction fine-tuning template for the image: http://farm6.staticflickr.com/5106/5670500150_e035dd2d30_z.jpg
+<SYSTEM>You are an AI visual assistant.
+Answer all questions as you are seeing the media<SYSTEM><USER>Assume the given image is broken into a 3X3 grid the regions or tiles being named upper-left upper-center”, upper-right”, middle-left”, middle-center”, middle-right”, bottom-left”, bottom-center”, bottom-right”.
+Rank these regions or tiles based on their saliency, give me the line separated ranking of all regions in decreasing order.
+<ASSISTANT>
+middle-right
+bottom-center
+bottom-right
+upper-center
+upper-right
+middle-center
+upper-left
+middle-left
+bottom-left
 The Volkswagen ad titled begins with a group of three women seated excitedly in a Volkswagen Jetta, the driver sporting a wide smile as she grips the steering wheel.
@@
 More such examples are given in Figs.
-4, 5, 7, and Figs.
-6, 8 for images and videos respectively.
+6, 7, 10, and Figs.
+8, 9 for images and videos respectively.
 In this captivating painting, a majestic tree takes center stage against a stunning backdrop.
@@
 Figure 6: Dense caption generated by Behavior-LLaVA for an artistic drawing of a tree.
-After collecting user behavior over image and video content, next, we design tasks to teach large vision and language models (VLMs) to simulate user behavior.
-For this, we use an instruction fine-tuning format.
-Given a video or an image and the other metadata like time of post and channel, we ask the model to simulate user behavior of likes and comments.
-See Fig 2, Listing 1 for examples.
-We choose LLaMA-Vid [52] as our base model to teach it the user behavior.
-We call the resultant model Behavior-LLaVA (Large Language and Vision Assistant) [54].
-Note that all the vision language models like LLaVA [54], MiniGPT-4 [91], Kosmos [32], BLIP [49], LLaMA-Vid [52] are trained on image-text instructions where text is some kind of annotations of the image, e.g., objects, object locations, captions, question-answers, etc.
-These instructions, since being sourced from either existing benchmark datasets such as MSCOCO [53] and Conceptual Captions [72] or being generated by stronger LLMs like GPT-4, require a significant amount of investment in terms of money for human annotation or generation through API.
-While building on top of their work, it is noteworthy that our model does not add any extra annotation or API cost to generate data.
-Rather, we rely on sampling data from digital analytics databases, which are anyway being collected for various reasons, and use them to improve VLM capabilities.
-Therefore, any potential improvements brought by integrating action data with VLMs do not require a significant capital investment.
 The image features a pair of Nike Air Force 1 sneakers, which are a classic and iconic design.
@@
 The red-colored text highlights the most important aspects of the video captured by Behavior-LLaVA, demonstrating an understanding of aesthetics, characters, world knowledge, emotion, and spatial relationships.
-We start with the trained LLaMA-Vid model.
-The LLaMA-Vid model uses two tokens to represent each video, which they call content and context tokens.
-While the context token encodes the overall image context based on user input, the content token encapsulates visual cues in each frame.
-For learning context tokens, the model uses attention queries that interact with previously generated image features in the designed attention module.
-To generate content tokens, the image features are average pooled.
-This dual-token strategy significantly reduces the number of tokens needed to represent videos, thus enabling the model to scale to longer (hour-long) videos.
-To better support hour-long videos, LLaMA-Vid was trained on a 9k movie-level conversation instruction set containing plot reasoning and detail understanding questions.
-Taking the base LLaMA-Vid model, we finetune it further on behavioral data.
-We teach the model a single instruction format given in Listing 1, where we teach it that given a media file (image or a video), it has to simulate the likes and comments on the media.
-We do this on a 730k instructions set created by us using Reddit and YouTube videos and images (§2).
-Next, we test the resultant model which we call Behavior-LLaVA, on a hierarchy of tasks in both zero-shot and finetuned settings.
-We include tasks based on low-level content understanding, like object and activity detection, and also on high-level understanding, like topic, sentiment, emotion, and persuasion classification.
-We also test the model on even higher-level tasks like memorability simulation and YouTube dislike/like simulation.
-We compare Behavior-LLaVA against the base model LLaMA-Vid and other state-of-the-art approaches reported on each test task (Tables 1-8).
-We make the following contributions with this work:
-1) We explore the idea of learning human behavior, resulting in better content understanding.
-We test this for action-level behavior data corresponding to receiver comments and likes.
-We collect a dataset consisting of 400k images and 330k videos along with their receiver comments and likes.
-Then, LLaMA-Vid is trained for the task of predicting receiver comments and likes given a media (a video or an image) (Listing 1).
-We show that using this simple task formulation and behavioral data collected in the wild, results in performance improvement over a hierarchy of tasks.
-We get improvements over the base LLaMA-Vid across 40 tasks over 23 benchmark datasets in both zero-shot and fine-tuned settings.
-We show this over low-level content understanding tasks like object and activity recognition and also over high-level tasks like topic and emotion detection.
-Further, interestingly, none of the tasks tested show a decrease in performance by training LLaMA-Vid on this additional task.
-Through this, we propose a scalable approach to increase the content understanding abilities of VLMs, requiring minimal cost and no architectural changes.
-2) To disentangle the effect of training LLaVA-Vid on the additional 730k video and image samples from the effect of training on behavior data over those samples, we train LLaVA-Vid on 730k samples using the original task of LLaVA-Vid.
-We call this model Ad-LLaVA.
-We show that Ad-LLaVA shows slight performance improvements over LLaVA-Vid; however, Behavior-LLaVA still performs better than both Ad-LLaVA and LLaMA-Vid, thus highlighting the importance of behavior data and instruction fine-tuning on behavior data.
-3) We also show an ablation of Behavior-LLaVA across different kinds of behavior.
-We try out the perception behavior of saliency prediction over images and two types of action-level behavior (comments and likes) over images and videos.
-We find that perception-level behavior does not result in significant performance improvements; however, action-level behavior shows improvements across all the tasks.
-We posit that one reason for this could be due to the scale for which action-level data is available (Table 9).
-While perception behavior is mostly collected in lab settings, action-level behavior data can be collected in a scalable manner automatically and cheaply.
-2 Methodology
-In this section, we introduce our approach to train Behavior-LLaVA.
-Since no publicly available corpus consists of behavior together with image and video content, we first introduce our instruction fine-tuning dataset, “Behavior-LLaVA Instruction Fine-Tuning dataset” (BLIFT).
-Next, we introduce our methodology to train Behavior-LLaVA.
-Finally, we report the results of testing Behavior-LLaVA’s capabilities on a hierarchy of tasks.
-The tasks cover low-level media understanding tasks like object and activity detection, high-level media understanding tasks like emotion, topic, and persuasion strategy classification.
-2.1 BLIFT Dataset
-Given the abundance of media and behavioral data and its accessibility, our data collection relies on two primary sources: Reddit and YouTube.
-These platforms share similarities in terms of hosting media content (images and videos) and user engagement metrics such as Reddit upvotes and comments, and YouTube likes, views, and comments.
-Here, we delineate the process involved in constructing the instruction fine-tuning dataset, which we term as the Behavior-LLaVA Instruction Fine-Tuning (BLIFT) dataset.
-2.1.1 Data from Reddit
-To collect a substantial corpus of diverse images and videos, we targeted two specific subreddits, namely r/pics and r/videos.
-Established over 15 years ago, these subreddits had a user base exceeding 20 million during the data collection period, with an average of over 5,000 online users concurrently.
-Notably, due to stringent content moderation guidelines [69, 2, 3, 4] and the exclusive focus of these subreddits on media content, they offer a rich variety of content devoid of thematic biases.
-Our data collection spans until January 2022, during which the activity on these subreddits witnessed a notable decline following several policy adjustments and user protests [29, 24].
-To ensure data quality and relevance, we executed a series of filtering steps on the posts and comments from these subreddits.
-Initially, we excluded posts predating February 2015 from r/pics, coinciding with the implementation of a rule requiring images without digital/overlay text [2, 3].
-This filtering step resulted in the exclusion of 3.1 million images and 2 million videos.
-Subsequently, considering the sustained popularity of both subreddits, with rankings within the top 20 since 2017 and consistent membership exceeding 20 million, we confined our dataset to posts from January 2018 onwards.
-This selection process yielded 1.4 million images and 1.1 million videos.
-Further refinement of the dataset involved removing posts and comments marked as NSFW, BOT-generated, or [deleted], along with eliminating duplicate images and videos.
-This curation step reduced the dataset to 876,000 images and 983,000 videos.
-To address redundancy in comments, we excluded those comprising fewer than three words and employed TF-IDF-based deduplication with a similarity threshold of 0.6, determined through manual observations.
-Following these steps, posts with fewer than two comments were filtered out, resulting in a dataset comprising 631,000 images and 397,000 videos.
-Additionally, videos exceeding a duration of 500 seconds were omitted, leaving 221,000 videos for analysis.
-Notably, images not directly hosted on Reddit were excluded due to scraping and copyright limitations.
-Similarly, for r/videos, only videos hosted on YouTube were considered.
-It is pertinent to mention that approximately 51% of YouTube videos collected during this period were either made private/unlisted or removed, resulting in 400,000 images and 80,000 videos, accompanied by 1.5 million and 312,000 comments, respectively.
-These comprehensive filtering steps ensured the construction of a diverse and relevant dataset for fine-tuning instruction-based models.
-Listing 1: Behavior Instruction fine-tuning template for the video: https://www.youtube.com/watch?v=BKPQkjRF4yY
-<SYSTEM>You are an AI visual assistant, you are given a detailed description of a media, followed by the actual media.
-Answer all questions as you are seeing the media<SYSTEM><USER>The video advertisement is titled Gatorade | Make Your Own Footsteps with Suni Lee for the brand Gatorade.
-The audio in the ad says ”[ASR HERE …]”.
-Analyze this video deeply, then write scene by scene description of the video and answer the following questions.
-What percentage of viewers would like this video, and what would be the top-5 popular comments on this video?.
-<video>.. </video>
-<USER><ASSISTANT>
-The scene by scene descriptions are:
-Scene 1: The scene shows a woman looking off into the distance with an orange line going around the outside of the screen.
-The foreground colors of the scene are Black,Mud_Green,Gray,Dark_Gray, and the background colors are Dark_Brown,Black,Dark_Gray.
-The dominant tone of the scene is neutral.
-This scene is categorized by the tags cosmetic, eyebrow, face, girl, ponytail, stand, string, woman.
-Scene 2: The scene shows a woman balancing on a skate board in a yard.
-The foreground colors of the scene are Black,Mud_Green,Dark_Gray,Olive, and the background colors are Black,Dark_Gray,Gray,Dark_Brown.
-The dominant tone of the scene is neutral.
-This scene is categorized by the tags athletic, balance, beam, car, girl, house exterior, hurdle, jog, legging, plank, rail, seesaw, woman, yard.
-Scene 3: The scene shows a girl jumping over a wooden ramp in the backyard.
-The foreground colors of the scene are Black,Dark_Gray,Gray,Dark_Blue, and the background colors are Dark_Brown,Dark_Blue,Purple,Dark_Pink,Brown.
-The dominant tone of the scene is neutral.
-This scene is categorized by the tags arm, backyard, balance, diving board, girl, house exterior, ledge, plank, purple, rail, shirt, short, stand, stretch, woman, yard.
-The video will be liked by 2.0%
-1. Wow. Love it. Shes such an inspiration to the the next generation as well as everyone.
-2. Inspiring and great story behind this commercial. Builds meaning and purpose in the hearts and mind of youth. Its been awhile since good meaningful ads have been made.
-3. Shes an inspiration to the world thanks to her my niece learning Gymnastics hopefully someday, she is an inspiration to other youngers as how suni is an inspiration to everyone
-4. Chills watching this. Such an inspiration
-5. Yooooo, this is straight up !
-<ASSISTANT>
-2.1.2 Data from YouTube
-Our data collection from YouTube begins with querying Wikidata [76] for YouTube IDs to compile a list of channels.
-Wikidata, derived from Wikipedia, provides a curated selection of renowned channels, automatically filtering out noisy videos commonly found in datasets collected from diverse sources like user-generated videos.
-This initial step yielded a dataset of 2.2 million videos spanning the period from 2018 to 2023, sourced from approximately 6,000 channels collected from Wikidata.
-To refine the dataset, manual filtering was employed to exclude certain categories deemed less relevant for our purposes.
-These categories included music and songs, gaming content, non-English videos, sports commentary, anime, memes, channels with disabled comments sections, and news-related content.
-Furthermore, videos exceeding a duration of 60 seconds were omitted, and only videos with a substantial viewership, defined as greater than 10,000 views, were retained.
-We observed that these videos usually have less noisy comments and likes.
-Subsequently, the top comments from each video, as ordered by YouTube (i.e., the most liked comments), were selected for inclusion in the dataset.
-To address redundancy in comments, a TF-IDF filter was applied with a threshold of 0.7, which proved effective in removing duplicate comments prevalent in YouTube data.
-Comments were further filtered to include only those with a minimum of four words and a maximum of 100 words, ensuring a balance between relevance and conciseness.
-Additionally, to mitigate the presence of NSFW content, a vocabulary specific to NSFW terms [1] was employed to filter out inappropriate posts.
-On average, we finally get 3.1 comments per video, providing a substantial corpus of user-generated content for analysis.
-After applying these filtering steps, the dataset was reduced to 250,000 videos, ensuring a curated and relevant collection for subsequent analysis and model training.
-2.2 Instruction Fine-Tuning LLaMA-Vid
-After collecting Reddit and YouTube media and user behavior, we formulate instruction fine-tuning tasks for training LLaMA-Vid.
-In the training instruction, given the media content and automatic speech recognition if available, we ask the model to simulate the scene-by-scene description and user likes/views and top-5 comments.
-This instruction training template is given in Listing 1.
-To generate the instruction data, first, frames are sampled using the 30-degree rule [73, 5, 26], then the scene-by-scene description are obtained by concatenating automatically generated captions and tags from LLaVA-13B [54], colors and tone through Qin et al.
-[67].
-This instruction format keeps the instructions similar to the instruction format for other VLMs like LLaVA [54], MiniGPT-4 [91], BLIP [48], LLaMA-Vid [52], etc., while additionally teaching the model to learn behavior.
-We keep the instruction fine-tuning template similar for both YouTube and Reddit.
-The complete instruction is given in Listing 1.
-To disentangle the effect of training on additional data samples from the effect of training on behavioral data, we train LLaMA-Vid on BLIFT with the video and image verbalization and do not include receiver behavior of comments and likes.
-Then, the overall instruction template consists of scene-by-scene automatically generated verbalization similar to Listing 1 without the likes and comment simulation.
-We call the LLaMA-Vid fine-tuned on this data, Ad-LLaVA.
-We compare Behavior-LLaVA with Ad-LLaVA and LLaMA-Vid along with other state-of-the-art literature benchmarks on various tasks (Tables 1-8).
-Listing 2: Perceptual Signal Instruction fine-tuning template for the image: http://farm6.staticflickr.com/5106/5670500150_e035dd2d30_z.jpg
-<SYSTEM>You are an AI visual assistant.
-Answer all questions as you are seeing the media<SYSTEM><USER>The objects in this image in no particular order are car, dog, frisbee.
-Give me the order of saliency of these objects, start with the most salient object and end with the least salient object, each in a separate line.
-Give me the objects only and nothing else.
-<image>
-<ASSISTANT>
-dog
-frisbee
-car
-<ASSISTANT>
-As an ablation experiment, we also try teaching the Behavior-LLaVA perceptual signals.
-For this, we take the largest perception signal dataset in the literature - Salicon10k [36].
-It consists of 10,000 MS COCO images [18] with free-viewing eye gaze data collected through a novel mouse-based interface.
-The dataset has been widely used in many studies.
-We formulate two tasks using this data, (1) Salicon [Object]: estimating the saliency over the objects in the image and (2) Salicon [Region]: estimating the saliency over a region, where the regions tiles formed by breaking the image into a 3x3 grid.
-For both tasks we try to model two objectives, ranking and predicting, we found ranking to be much more effective.
-The instruction is given in Listings 2 and 3.
-Listing 3: Perceptual Signal Instruction fine-tuning template for the image: http://farm6.staticflickr.com/5106/5670500150_e035dd2d30_z.jpg
-<SYSTEM>You are an AI visual assistant.
-Answer all questions as you are seeing the media<SYSTEM><USER>Assume the given image is broken into a 3X3 grid the regions or tiles being named upper-left upper-center”, upper-right”, middle-left”, middle-center”, middle-right”, bottom-left”, bottom-center”, bottom-right”.
-Rank these regions or tiles based on their saliency, give me the line separated ranking of all regions in decreasing order.
-<ASSISTANT>
-middle-right
-bottom-center
-bottom-right
-upper-center
-upper-right
-middle-center
-upper-left
-middle-left
-bottom-left
-We take the pretrained checkpoints of the LLaMA-VID 13B [51] model after the modality-alignment stage.
-We combine our instruction pairs from BLIFT with the original instruction tuning dataset consisting of 40K text conversations from ShareGPT, 625K single or multi-turn visual QA pairs, and 98K video QA pairs, all the modules except the Visual Encoder are kept frozen.
-We train the LLaMA-Vid checkpoints with their original sft mix along with BLIFT.
-We ablate different sampling ratios and found 1:1 to be empirically performing the best.
-We train the model for 2.2 epochs, keeping track of the 0-shot evaluation metrics and perplexity on comments in the eval set.
-For the best checkpoint, the perplexity on comments reduces from 6.22 to 3.05, and the $R^{2}$ on likes/views goes from -5.1 to 0.45.
-3 Results and Discussion
-In the experimental results, we aim to showcase the diverse and emergent capabilities of our Behavior-LLaVA model through quantitative numbers on various tasks and qualitative examples.
-These abilities include generating detailed image and video descriptions, emotion and sentiment analysis, question answering, video understanding tasks like scene and action detection.
-Additionally, we present the ability of Behavior-LLaVA to transfer learn on other behaviors like memorability of a video - both short-term and long-term.
-Model scene way_speaking relationship like_ratio view_count director genre writer year Video4096-GPT-3.5 generated story + Flan-t5-xxl [13] 60.2 39.07 64.1 0.061 12.84 69.9 58.1 52.4 75.6 Video4096-GPT-3.5 generated story + GPT-3.5 classifier [13] 54.54 32.95 68.42 0.031 12.69 75.26 50.84 32.16 75.96 LLaMA-Vid + GPT-3.5 Generated Story 58.12 35.5 60.6 0.314 10.34 65.34 49.77 34.23 72.12 Ad-LLaVA 59.05 37.07 61.2 0.319 10.37 66.84 55.13 35.33 77.34 Behavior LLaVA + GPT-3.5 Generated Story 66.43 41.03 64.21 0.17 5.12 71.12 63.45 39.4 79.3 Improvement of Behavior LLaVA over LLaMA-Vid 10.48% 15.58% 9.62% 45.86% 50.48% 8.85% 27.49% 15.1% 9.96%
-Table 1: Comparison of various models on the Long Video Understanding benchmark [79] consisting of 9 VQA tasks.
-We see that Behavior-LLaVA improves on LLaMA-Vid on 9/9 tasks with an average improvement of 21.49%.
-Further, it outperforms the state-of-the-art in 5/9 tasks.
-Model Scene Object Action Event Attribute Concept Overall 0-shot LLaMA-VID 47.25 65.26 78.12 28.03 42.33 50.03 51.83 Ad-LLaVA 49.10 65.35 77.45 31.45 43.33 50.70 52.91 Behavior-LLaVA 52.03 65.33 77.95 32.66 45.67 51.20 54.14 Improvement of Behavior-LLaVA over LLaMA-Vid 10.12% 0.11% -0.22% 16.52% 7.89% 2.34% 4.46% 0-shot w/ story Video4096 - GPT-3.5 generated story + Flan-t5-xxl classifier [13] 59.66 98.89 98.96 38.42 67.76 86.99 75.12 Video4096 - GPT-3.5 generated story + GPT-3.5 classifier [13] 60.2 99.16 98.72 40.79 67.17 88.6 75.77 LLaMA-VID + Generated Story 60.3 99.92 99.01 39.33 66.66 87.33 75.425 Behavior-LLaVA + GPT3.5 generated story 60.4 99.89 98.23 40.97 67.23 88.33 75.84 Improvement of Behavior-LLaVA over LLaMA-Vid 0.16% -0.03% -0.78% 4.17% 0.85% 1.14% 0.55%
-Table 2: Comparison of various models on the Holistic Video Understanding benchmark [23] consisting of 7 VQA tasks.
-We see that Behavior-LLaVA improves on LLaMA-Vid on 6/7 tasks with an average improvement of 5.88%.
-Further, it performs equivalent to the state-of-the-art in 6/7 tasks.
-State of the art is achieved by generating a story and asking Behavior-LLaVA to answer questions based on the generated story.
-3.1 Evaluation
-To test the effectiveness of Behavior-LLaVA, we conduct experiments involving 40 distinct tasks across 23 benchmark datasets.
-The diversity of tasks and datasets allows us to evaluate the performance and capabilities of Behavior-LLaVA thoroughly.
-Each of them is covered briefly next:
-Method LLM MSVD-QA MSRVTT-QA ActivityNet-QA Acc Score Acc Score Acc Score FrozenBiLM [84] DeBERTa-V2 32.2 16.8 24.7 VideoLLaMA [87] Vicuna-7B 51.6 2.5 29.6 1.8 12.4 1.1 LLaMA-Adapter [88] LLaMA-7B 54.9 3.1 43.8 2.7 34.2 2.7 VideoChat [50] Vicuna-7B 56.3 2.8 45.0 2.5 26.5 2.2 Video-ChatGPT [56] Vicuna-7B 64.9 3.3 49.3 2.8 35.2 2.7 BT-Adapter [55] Vicuna-7B 67.5 3.7 57.0 3.2 45.7 3.2 LLaMA-VID Vicuna-7B 69.7 3.7 57.7 3.2 47.4 3.3 LLaMA-VID Vicuna-13B 70.0 3.7 58.9 3.3 47.5 3.3 Ad-LLaVA Vicuna-13B 70.0 3.7 59.0 3.3 47.4 3.3 Behaviour-LLaVA Vicuna-13B 70.1 3.7 59.2 3.4 47.5 3.3 Improvement of Behavior-LLaVA over LLaMA-Vid 0.14% 0% 0.5% 3% 0% 0%
-Table 3: Comparison of various models on three conventional video question answering benchmarks consisting of question answers related to action understanding.
-The main goal of comparing on this benchmark is to show that Behavior-LLaVA does not perform worse on low-level understanding tasks like action recognition.
-We see that Behavior-LLaVA marginally improves on LLaMA-Vid on 2/3 benchmarks.
-Further, it performs equivalent to the state-of-the-art in 3/3 benchmarks.
-  1. 1.
-    Visual Question Answering (VQA): We evaluate the performance of visual question answering on the following benchmark datasets:
-    • The Long-Video Understanding (LVU) benchmark by Wu and Krahenbuhl [79] comprises nine distinct tasks aimed at assessing long video comprehension, incorporating over 1000 hours of video content.
-These tasks encompass diverse aspects such as content understanding (including ‘relationship’, ‘speaking style’, ‘scene/place’), prediction of user engagement (‘YouTube like ratio’, ‘YouTube popularity’), and forecasting movie metadata (‘director’, ‘genre’, ‘writer’, ‘movie release year’).
-Evaluation of LVU tasks by Wu and Krahenbuhl [79] utilizes top-1 classification accuracy for content understanding and metadata prediction tasks, while Mean Squared Error (MSE) is employed for user engagement prediction tasks.
-    • The Holistic Video Understanding (HVU) dataset, as introduced by Diba et al.
-[23], stands as the largest dataset for long video comprehension, comprising 572,000 samples.
-Encompassing a broad spectrum of semantic elements within videos, HVU tasks involve the classification of scenes, objects, actions, events, attributes, and concepts.
-Performance evaluation on HVU tasks, akin to the original work, is conducted using the mean average precision (mAP) metric on the validation set.
-    • MSVD-QA and MSRVTT-QA: These datasets are based on Microsoft Research Video Description [17] and MSR-VTT corpora [82] and are extensively used in many video captioning and question-answering experiments.
-The MSVD-QA dataset has a total number of 1,970 video clips and 50,505 question-answer pairs.
-The MSRVTT-QA dataset contains 10K video clips and 243k question-answer pairs.
-    • ActivityNet-QA [16] is a benchmark primarily for human activity understanding containing 849 hours of video, including 28,000 action instances.
-  2. 2.
-    Video and Image Understanding Benchmarks: We use a wide variety of tasks to evaluate video and image understanding: topic, emotion, and persuasion strategy classification, action and reason retrieval and generation, and emotions.
-We briefly introduce the benchmarks:
-    • The advertisements dataset by Hussain et al.
-[33] contains 3,477 video advertisements and the corresponding annotations for emotion and topic tags and action-reason statements for each video.
-There are a total of 38 topics and 30 unique emotion tags per video.
-Further, we have 5 action-reason statements for each video for the action-reason generation task.
-    • Persuasion strategy dataset [13] is a dataset consisting of 1002 video advertisements from popular brands and their persuasion strategy labels like social identity, anchoring and comparison, reciprocity, foot-in-the-door, etc.
-      Training Model Topic Sentiment Persuasion Action Reason Clubbed All labels Random Random 2.63 3.37 14.3 8.37 3.34 3.33 Zero-shot VideoChat [50] 9.07 3.09 5.1 10.28 - - Video4096 - GPT-3.5 Generated Story + GPT-3.5 Classifier [13] 51.6 11.68 79.69 35.02 66.27 59.59 Video4096 - GPT-3.5 Generated Story + Flan-t5-xxl Classifier [13] 60.5 10.8 79.10 33.41 79.22 81.72 Video4096 - GPT-3.5 Generated Story + Vicuna Classifier [13] 22.92 10.8 67.35 29.6 21.39 20.89 Video4096 - Vicuna Generated Story + GPT-3.5 Classifier [13] 46.7 5.9 80.33 27.54 61.88 55.44 Video4096 - Vicuna Generated Story + Flan-t5-xxl Classifier [13] 57.38 9.8 76.60 30.11 77.38 80.66 Video4096 - Vicuna Generated Story + Vicuna Classifier [13] 11.75 10.5 68.13 26.59 20.72 21.00 LCBM [39] 42.17 7.08 58.83 32.83 39.55 27.91 LLaMA-VID w/ only video 10.11 3.42 5.75 12.32 29.61 24.11 LLaMA-VID w/ video + GPT-3.5 Story 42.72 11.05 64.02 32.07 37.76 42.33 Behavior-LLaVA w/ only video 22.65 11.13 60.04 13.39 42.66 33.33 Behavior-LLaVA w/ video + verbalization 46.34 11.7 64.13 33.33 52.06 52.03 Ad-LLaVA w/ video + GPT-3.5 story 51.16 11.33 68.03 33.11 43.26 51.45 Behavior-LLaVA w/ video + GPT-3.5 story 60.09 12.84 79.94 36.12 67.10 79.18 Improvement of Behavior-LLaVA over LLaMA-Vid 40.66% 16.2% 24.86% 12.62% 77.7% 87.05% Finetuned VideoMAE [75] 24.72 29.72 85.55 11.17 - - Hussain et al.
-[33] 35.1 32.8 - - 48.45 - Intern-Video [78] 57.47 36.08 86.59 5.47 6.8 - Video4096- Generated Story + Roberta Classifier [13] 71.3 33.02 84.20 64.67 42.96 39.09 LLaMA-VID w/ video + verbalization 59.13 32.11 79.15 50.93 50.32 30.13 LLaMA-VID w/ video + GPT-3.5 Story 63.11 35.01 84.15 55.01 57.11 45.73 Behavior-LLaVA w/ only video 58.03 22.72 84.41 26.23 59.33 51.45 Behavior-LLaVA w/ video + verbalization 68.32 33.92 85.93 64.72 70.89 75.34 Ad-LLaVA w/ video + GPT-3.5 story 66.34 36.24 84.09 58.31 68.15 78.15 Behavior-LLaVA w/ video + GPT-3.5 story 71.2 39.55 86.17 65.03 80.44 81.67 Improvement of Behavior-LLaVA over LLaMA-Vid 12.82% 12.97% 2.4% 18.21% 40.85% 78.59%
-      Table 4: Comparison of various models on two video understanding benchmarks [33, 44] consisting of 5 tasks related to video advertisements understanding.
-The main goal of comparing on these benchmarks is to demonstrate Behavior-LLaVA’s understanding of complex videos.
-We see that Behavior-LLaVA improves on LLaMA-Vid on 5/5 tasks with an average improvement score of 43.18% in zero-shot and 27.64% in fine-tuned settings.
-Further, it outperforms the current state-of-the-art on 3/5 tasks in zero-shot and 5/5 in fine-tuned settings.
-    • VideoEmotion-8 [7] dataset comprises 1,101 user-generated videos sourced from YouTube and Flickr, each containing a minimum of 100 videos per emotional category, as per Plutchik Wheel’s emotion model.
-    • Ekman-6 [81] dataset is compiled from social websites, with each of its 1,637 videos labeled with a single emotion category based on Ekman’s psychological research.
-    • CAER [47] dataset, sourced from TV shows, consists of 13,201 clips with an average sequence length of 90, each manually labeled with six basic emotions, aligning with the Ekman-6 dataset.
-    • IAPSa [60] is a subset of IAPS, following the Mikels model with eight emotion categories such as amusement, awe, contentment, excitement, anger, disgust, fear, and sadness.
-It consists of 395 affective images, marking the first visual emotion dataset with discrete categories.
-    • Emotion6 [64] features 1,980 images sourced from Flickr, each labeled by 15 annotators according to the Ekman model, covering six emotion categories: happiness, anger, disgust, fear, sadness, and surprise.
-    • EmoSet [86] encompasses a total of 3.3 million images, including 118,102 from social networks and artistic sources, evenly distributed across various emotion categories.
-Based on the Mikels model, EmoSet is categorized into eight emotion categories.
-    • Abstract [57] exclusively consists of color and texture combinations without recognizable objects.
-Differing from the IAPS dataset where emotions often stem from identifiable objects, the abstract paintings dataset was peer-rated via a web survey, with each image rated approximately 14 times.
-It comprises 228 images spanning eight categories similar to those in IAPS.
-  3. 3.
-    Image Dense Captioning: Literature image captioning datasets such as MS-COCO [18] reduce the inherently rich information and fine-grained semantics to simplistic captions, with very brief statements often focussing only on salient objects.
-This leads to a significant reduction in information content.
-Behavior data such as user comments help a model learn much more information such as object and material properties, world knowledge, emotion, character understanding, spatial relationships, aesthetics, etc.
-(see Fig.
-1), enhancing the model’s captioning capability.
-Therefore, we design a captioning task to test this capability and compare it with respect to LLaMA-Vid and LLaVA-34B (a 2.5x larger model).
-Since we do not have ground truth for this task, following the LLM-as-a-judge paradigm, we use GPT-4V as the judge for all the models.
-GPT-4V is asked to evaluate the dense captions on three metrics: Correctness (Listing 4) evaluating the factuality and model hallucinations, Detail (Listing 5) evaluating the number and depth of details captured by the generated captions, and Quality (Listing 6) measuring the subjective quality of the concepts chosen to be highlighted by the captioning model and the arrangement, coherence, and the linking of various concepts.
-  4. 4.
-    Image and Video Memorability Simulation: Behavior-LLaVA is trained on behavior along with the media.
-To check if training on behavior helps in solving other behavior tasks [39], we test it over image and video memorability simulation.
-For this, we select seven benchmark datasets covering long-term and short-term memorability over images and videos.
-    • Memento10k[61] is a short-term video memorability dataset comprising 10,000 video clips, with 900,000 human memory annotations recorded at various delay intervals.
-The video clips were, on average, 3s long.
-    • VideoMem [20] is comprised of 10,000 soundless videos, each lasting 7 seconds, accompanied by memorability scores.
-Memorability was measured twice: first, shortly after viewing and again 24-72 hours later to capture both short-term and long-term memorability effects.
-    • LaMem [40] dataset is a short-term image memorability dataset comprising of 60000 images.
-The dataset contains scene-centric images, object-centric images and other types such as images of art, images evoking certain emotions, and other user-generated images.
-    • SUN [35] dataset is a short-term image memorability dataset comprising $2$ images that were sourced from the SUN database.
-      Training Dataset Video Emotion-8 CAER Ekman-6 Random Random 12.5 14.28 16.67 0-Shot LLaMA-Vid 29.7 27.2 37.33 Behavior-LLaVA 41.35 51.0 49.33 Ad-LLaVA 29.8 27.3 37.66 Improvement of Behavior-LLaVA over LLaMA-Vid 39.22% 84.19% 32.14% Finetuned Zhao et al.
-[90] 54.5 78.3 55.3 Zhang et al.
-[89] 57.3 80.1 58.2 eMOTIONS [80] - - 53.12 Arevalo et al.
-[6] 53.7 77.3 54.2 Qiu et al.
-[68] 53.3 - 57.3 Xu et al.
-[81] 52.6 77.9 55.6 LLaMA-Vid 53.8 75.6 57.9 Ad-LLaVA 54.1 76.1 57.8 Behavior-LLaVA 56.9 79.3 58.4 Improvement of Behavior-LLaVA over LLaMA-Vid 5.76% 3.57% 0.86%
-      Table 5: Comparison of various models on three video emotion understanding benchmarks (Video Emotion8 [37], CAER [47], Ekman-6 [81]).
-The main goal of comparing on these benchmarks is to demonstrate Behavior-LLaVA’s understanding of complex tasks like video emotions of long-form videos.
-We see that Behavior-LLaVA improves on LLaMA-Vid on 3/3 benchmarks with an average improvement score of 51.85% in zero-shot and 3.39% in fine-tuned settings.
-Further, it outperforms the current state-of-the-art on 3/3 benchmarks in zero-shot and 1/3 in fine-tuned settings.
-    • MemCat [27] dataset is a short-term image memorability dataset comprising 10000 images.
-It consists of five broader memorability-relevant semantic categories (animal, sports, food, landscapes, vehicles), with 2K exemplars each, further divided into different subcategories (e.g., bear, pigeon, cat, etc.
-for animal).
-The images were sourced from existing image sets: ImageNet, COCO, Open Images Dataset, and SUN.
-    • MediaEval [42] utilized publicly available links to short-form video clips, each averaging 6 seconds in duration, with both short-term and long-term memorability scores.
-Short-term memorability evaluations were conducted on videos viewed within the preceding few minutes, while long-term memorability assessments were based on videos viewed within the previous 24 to 72 hours.
-    • LAMBDA [34] is a long-term memorability consisting of 2205 video ads collected over 1749 participants covering 276 brands.
-The average video length is 33 seconds and the videos are highly complex, consisting of audio, logos, fast-moving scenes, emotions, etc.
-For Tables 1, 2, and 4, we follow the evaluation protocol of Video-4096 [13], for Table 3 we follow the evaluation protocol of LLaVA and LLaMA-VID.
-For Tables 5, 6, and 8, for 0-shot evaluation results, we use the logits of the next token from the given task vocabulary.
-For Table 8, we use the evaluation protocol of SI et al.
-[73].
 In a heart-pounding and visually stunning trailer for Red Dead Redemption, we are thrust into the gritty world of the American frontier.
@@
 The original video is posted at URL: https://www.youtube.com/watch?v=eaW0tYpxyp0.
-The composition of the image presents a compelling narrative of a soldier immersed in a war-torn landscape.
-Positioned amidst a backdrop of explosive chaos and dense foliage, the soldier, adorned in traditional military gear, gazes directly at the viewer with a resolute demeanor.
-The soldier’s face is camouflaged, blending seamlessly with the olive-green helmet—a hallmark of battlefield attire.
-Surrounding this central figure, a tableau of activity unfolds: a tank looms in the background, its barrel skyward, suggesting recent action.
-Other soldiers, alert and vigilant, navigate the dense jungle terrain, underscoring the high stakes of the conflict.
-The foliage, lush yet foreboding, heightens the palpable tension of the scene.
-Foregrounded by the silhouette of another soldier’s helmeted head, the viewer is drawn into the heart of the action, evoking a sense of shared experience amid the perils of warfare.
-The realism of the depiction accentuates the emotional weight of the moment, capturing the essence of human resilience amidst the ravages of battle.
-Overall, the image transcends mere representation, offering a poignant reflection on the individual’s role in the broader narrative of war—imbued with suspense, authenticity, and a profound exploration of the human condition within conflict.
-Figure 7: Dense caption generated by Behavior-LLaVA for a painting of a soldier. The model captures many qualitative aspects that are usually missed in common captioning tasks.
 The scene is set at a packed stadium buzzing with excitement, the air thick with anticipation as Argentina and France face off in what promises to be an epic World Cup final.
@@
 The original video is posted at URL: https://www.youtube.com/watch?v=zhEWqfP6V_w.
-Training Models IAPSa-8 Abstract Emotion6 Emoset Random Random 12.5 12.5 16.67 12.5 0-shot GPT4-V 83.33 71.12 65.47 79.16 LLaMA-Vid 43.41 43.24 40.37 45.23 Ad-LLaVA 43.22 43.01 43.21 44.38 Behavior-LLaVA 57.97 64.21 49.71 50.38 Improvement of Behavior-LLaVA over LLaMA-Vid 33.54% 48.5% 23.14% 11.39% Finetuned MIDAN [83] 85.96 78.34 61.66 75.75 Stimuli-aware [85] - - 61.62 78.40 LLaMA-VID finetuned 84.93 71.23 62.87 80.31 Ad-LLaVA finetuned 85.13 71.16 62.66 79.88 Behavior-LLaVA finetuned 87.36 81.41 72.31 83.21 Improvement of Behavior-LLaVA over LLaMA-Vid 2.86% 14.29% 15.02% 3.61%
-Table 6: Comparison of various models on four image emotion understanding benchmarks (IAPSa-8 [60] Abstract [57], Emotion6 [64], Emoset [86]).
+The composition of the image presents a compelling narrative of a soldier immersed in a war-torn landscape.
+Positioned amidst a backdrop of explosive chaos and dense foliage, the soldier, adorned in traditional military gear, gazes directly at the viewer with a resolute demeanor.
+The soldier’s face is camouflaged, blending seamlessly with the olive-green helmet—a hallmark of battlefield attire.
+Surrounding this central figure, a tableau of activity unfolds: a tank looms in the background, its barrel skyward, suggesting recent action.
+Other soldiers, alert and vigilant, navigate the dense jungle terrain, underscoring the high stakes of the conflict.
+The foliage, lush yet foreboding, heightens the palpable tension of the scene.
+Foregrounded by the silhouette of another soldier’s helmeted head, the viewer is drawn into the heart of the action, evoking a sense of shared experience amid the perils of warfare.
+The realism of the depiction accentuates the emotional weight of the moment, capturing the essence of human resilience amidst the ravages of battle.
+Overall, the image transcends mere representation, offering a poignant reflection on the individual’s role in the broader narrative of war—imbued with suspense, authenticity, and a profound exploration of the human condition within conflict.
+Figure 10: Dense caption generated by Behavior-LLaVA for a painting of a soldier. The model captures many qualitative aspects that are usually missed in common captioning tasks.
+Training Model Topic Sentiment Persuasion Action Reason Clubbed All labels Random Random 2.63 3.37 14.3 8.37 3.34 3.33 Zero-shot VideoChat (Li et al., 2023a) 9.07 3.09 5.1 10.28 - - Video4096 - GPT-3.5 Generated Story + GPT-3.5 Classifier (Bhattacharyya et al., 2023) 51.6 11.68 79.69 35.02 66.27 59.59 Video4096 - GPT-3.5 Generated Story + Flan-t5-xxl Classifier (Bhattacharyya et al., 2023) 60.5 10.8 79.10 33.41 79.22 81.72 Video4096 - GPT-3.5 Generated Story + Vicuna Classifier (Bhattacharyya et al., 2023) 22.92 10.8 67.35 29.6 21.39 20.89 Video4096 - Vicuna Generated Story + GPT-3.5 Classifier (Bhattacharyya et al., 2023) 46.7 5.9 80.33 27.54 61.88 55.44 Video4096 - Vicuna Generated Story + Flan-t5-xxl Classifier (Bhattacharyya et al., 2023) 57.38 9.8 76.60 30.11 77.38 80.66 Video4096 - Vicuna Generated Story + Vicuna Classifier (Bhattacharyya et al., 2023) 11.75 10.5 68.13 26.59 20.72 21.00 LCBM (Khandelwal et al., 2024) 42.17 7.08 58.83 32.83 39.55 27.91 LLaMA-VID w/ only video 10.11 3.42 5.75 12.32 29.61 24.11 LLaMA-VID w/ video + GPT-3.5 Story 42.72 11.05 64.02 32.07 37.76 42.33 Behavior-LLaVA w/ only video 22.65 11.13 60.04 13.39 42.66 33.33 Behavior-LLaVA w/ video + verbalization 46.34 11.7 64.13 33.33 52.06 52.03 Ad-LLaVA w/ video + GPT-3.5 story 51.16 11.33 68.03 33.11 43.26 51.45 Behavior-LLaVA w/ video + GPT-3.5 story 60.09 12.84 79.94 36.12 67.10 79.18 Improvement of Behavior-LLaVA over LLaMA-Vid 40.66% 16.2% 24.86% 12.62% 77.7% 87.05% Finetuned VideoMAE (Tong et al., 2022) 24.72 29.72 85.55 11.17 - - Hussain et al.
+(2017) 35.1 32.8 - - 48.45 - Intern-Video (Wang et al., 2022) 57.47 36.08 86.59 5.47 6.8 - Video4096- Generated Story + Roberta Classifier (Bhattacharyya et al., 2023) 71.3 33.02 84.20 64.67 42.96 39.09 LLaMA-VID w/ video + verbalization 59.13 32.11 79.15 50.93 50.32 30.13 LLaMA-VID w/ video + GPT-3.5 Story 63.11 35.01 84.15 55.01 57.11 45.73 Behavior-LLaVA w/ only video 58.03 22.72 84.41 26.23 59.33 51.45 Behavior-LLaVA w/ video + verbalization 68.32 33.92 85.93 64.72 70.89 75.34 Ad-LLaVA w/ video + GPT-3.5 story 66.34 36.24 84.09 58.31 68.15 78.15 Behavior-LLaVA w/ video + GPT-3.5 story 71.2 39.55 86.17 65.03 80.44 81.67 Improvement of Behavior-LLaVA over LLaMA-Vid 12.82% 12.97% 2.4% 18.21% 40.85% 78.59%
+Table 6: Comparison of various models on two video understanding benchmarks (Hussain et al., 2017; Kumar et al., 2023) consisting of 5 tasks related to video advertisements understanding.
+The main goal of comparing on these benchmarks is to demonstrate Behavior-LLaVA’s understanding of complex videos.
+We see that Behavior-LLaVA improves on LLaMA-Vid on 5/5 tasks with an average improvement score of 43.18% in zero-shot and 27.64% in fine-tuned settings.
+Further, it outperforms the current state-of-the-art on 3/5 tasks in zero-shot and 5/5 in fine-tuned settings.
+Method LLM MSVD-QA MSRVTT-QA ActivityNet-QA Acc Score Acc Score Acc Score FrozenBiLM (Yang et al., 2022) DeBERTa-V2 32.2 16.8 24.7 VideoLLaMA (Zhang et al., 2023a) Vicuna-7B 51.6 2.5 29.6 1.8 12.4 1.1 LLaMA-Adapter (Zhang et al., 2023b) LLaMA-7B 54.9 3.1 43.8 2.7 34.2 2.7 VideoChat (Li et al., 2023a) Vicuna-7B 56.3 2.8 45.0 2.5 26.5 2.2 Video-ChatGPT (Maaz et al., 2023) Vicuna-7B 64.9 3.3 49.3 2.8 35.2 2.7 BT-Adapter (Liu et al., 2023b) Vicuna-7B 67.5 3.7 57.0 3.2 45.7 3.2 LLaMA-VID Vicuna-7B 69.7 3.7 57.7 3.2 47.4 3.3 LLaMA-VID Vicuna-13B 70.0 3.7 58.9 3.3 47.5 3.3 Ad-LLaVA Vicuna-13B 70.0 3.7 59.0 3.3 47.4 3.3 Behaviour-LLaVA Vicuna-13B 70.1 3.7 59.2 3.4 47.5 3.3 Improvement of Behavior-LLaVA over LLaMA-Vid 0.14% 0% 0.5% 3% 0% 0%
+Table 7: Comparison of various models on three conventional video question answering benchmarks consisting of question answers related to action understanding.
+The main goal of comparing on this benchmark is to show that Behavior-LLaVA does not perform worse on low-level understanding tasks like action recognition.
+We see that Behavior-LLaVA marginally improves on LLaMA-Vid on 2/3 benchmarks.
+Further, it performs equivalent to the state-of-the-art in 3/3 benchmarks.
+Training Models IAPSa-8 Abstract Emotion6 Emoset Random Random 12.5 12.5 16.67 12.5 0-shot GPT4-V 83.33 71.12 65.47 79.16 LLaMA-Vid 43.41 43.24 40.37 45.23 Ad-LLaVA 43.22 43.01 43.21 44.38 Behavior-LLaVA 57.97 64.21 49.71 50.38 Improvement of Behavior-LLaVA over LLaMA-Vid 33.54% 48.5% 23.14% 11.39% Finetuned MIDAN (Xu et al., 2022) 85.96 78.34 61.66 75.75 Stimuli-aware (Yang et al., 2021) - - 61.62 78.40 LLaMA-VID finetuned 84.93 71.23 62.87 80.31 Ad-LLaVA finetuned 85.13 71.16 62.66 79.88 Behavior-LLaVA finetuned 87.36 81.41 72.31 83.21 Improvement of Behavior-LLaVA over LLaMA-Vid 2.86% 14.29% 15.02% 3.61%
+Table 8: Comparison of various models on four image emotion understanding benchmarks (IAPSa-8 (Mikels et al., 2005) Abstract (Machajdik & Hanbury, 2010), Emotion6 (Peng et al., 2015), Emoset (Yang et al., 2023)).
 The main goal of comparing on these benchmarks is to demonstrate Behavior-LLaVA’s understanding of complex tasks like image emotions.
@@
 On the aspects of detail and quality, it even outperforms the much larger model of LLaVA-1.6 (34B).
-Training Models Image Datasets Video Datasets Lamem Memcat SUN Memento10k VideoMem MediaEval LAMBDA Human Consistency 0.68 0.78 0.75 0.73 0.61 - 0.61 Finetuned 10-shot in-context learning GPT-3.5 0.29 0.18 0.15 0.07 0.06 0.06 0.06 ViTMem [28] 0.71 0.65 0.63 0.56 0.51 - 0.08 Henry trained with 25% data [34] 0.56 0.64 0.59 0.62 0.49 0.32 0.28 Henry trained with 50% data [34] 0.65 0.68 0.67 0.69 0.55 0.44 0.40 Henry trained with 75% data [34] 0.71 0.75 0.73 0.74 0.62 0.49 0.47 Henry trained on all (combined) datasets [34] 0.72 0.79 0.76 0.72 0.60 0.48 0.52 Ad-LLaVA trained with 50% data 0.67 0.65 0.61 0.69 0.56 0.43 0.47 Behaviour LLaVA trained with 25% data 0.67 0.72 0.69 0.68 0.53 0.44 0.50 Behaviour LLaVA trained with 50% data 0.72 0.77 0.73 0.71 0.59 0.46 0.51 Behaviour LLaVA trained with 75% data 0.73 0.77 0.74 0.70 0.60 0.47 0.50 Behavior-LLaVA trained on all datasets 0.73 0.78 0.74 0.71 0.60 0.47 0.52 Improvement of Behavior-LLaVA over LLaMA-Vid (25% data) 19.64% 12.5% 16.95% 9.68% 8.16% 37.5% 78.57% Improvement of Behavior-LLaVA over LLaMA-Vid (50% data) 10.77% 13.26% 8.96% 2.90% 7.27% 4.54% 27.5% 0-shot LLaMA-Vid 0.13 0.11 0.05 0.03 0.05 0.02 0.05 Ad-LLaVA 0.14 0.13 0.06 0.06 0.07 0.04 0.13 Behavior-LLaVA 0.21 0.17 0.13 0.12 0.08 0.07 0.16 Improvement of Behavior-LLaVA over LLaMA-Vid 61.5% 54.5% 160% 300% 160% 350% 219%
-Table 8: Comparison of various models on seven video and image memorability benchmarks (Memento10k [61], VideoMem [20], LaMem [40], SUN [35], MemCat [27], MediaEval [42], LAMBDA [34]).
-The main goal of comparing on these benchmarks is to demonstrate Behavior-LLaVA’s understanding of complex and high-level tasks like memorability simulation.
-We see that Behavior-LLaVA improves on LLaMA-Vid on 7/7 benchmarks with an average improvement score of 186.4% in zero-shot and 39% in fine-tuned settings after seeing 25% train data.
-Further, it performs similarly to the current state-of-the-art on 7/7 benchmarks in the fine-tuned settings while still seeing only 25% data.
-Task MSRVTT-QA Emoset Comments likes/views Base 58.9 45.23 6.22 -0.1 Salicon [Region] 55.6 47.25 6.25 -0.07 Salicon [Object] 57.9 48.0 6.12 0.05 BLIFT[Likes/Views] 58.2 47.12 6.15 0.38 BLIFT[Comments] 59.0 49.58 3.02 0.19 BLIFT 59.2 50.38 3.05 0.40
+Model Scene Object Action Event Attribute Concept Overall 0-shot LLaMA-VID 47.25 65.26 78.12 28.03 42.33 50.03 51.83 Ad-LLaVA 49.10 65.35 77.45 31.45 43.33 50.70 52.91 Behavior-LLaVA 52.03 65.33 77.95 32.66 45.67 51.20 54.14 Improvement of Behavior-LLaVA over LLaMA-Vid 10.12% 0.11% -0.22% 16.52% 7.89% 2.34% 4.46% 0-shot w/ story Video4096 - GPT-3.5 generated story + Flan-t5-xxl classifier 59.66 98.89 98.96 38.42 67.76 86.99 75.12 Video4096 - GPT-3.5 generated story + GPT-3.5 classifier 60.2 99.16 98.72 40.79 67.17 88.6 75.77 LLaMA-VID + Generated Story 60.3 99.92 99.01 39.33 66.66 87.33 75.425 Behavior-LLaVA + GPT3.5 generated story 60.4 99.89 98.23 40.97 67.23 88.33 75.84 Improvement of Behavior-LLaVA over LLaMA-Vid 0.16% -0.03% -0.78% 4.17% 0.85% 1.14% 0.55%
+Table 10: Comparison of various models on the Holistic Video Understanding benchmark (Diba et al., 2020) consisting of 7 VQA tasks.
+We see that Behavior-LLaVA improves on LLaMA-Vid on 6/7 tasks with an average improvement of 5.88%.
+Further, it performs equivalent to the state-of-the-art in 6/7 tasks.
+State of the art is achieved by generating a story and asking Behavior-LLaVA to answer questions based on the generated story.
+Task MSRVTT-QA CAER Emoset Comments likes/views Base 58.9 75.6 45.23 6.22 -0.1 Salicon [Region] 55.6 75.8 47.25 6.25 -0.07 Salicon [Object] 57.9 76.4 48.0 6.12 0.05 BLIFT[Likes/Views] 58.2 76.2 47.12 6.15 0.38 BLIFT[Titles] 58.4 78.1 48.12 5.09 0.13 BLIFT[Comments] 59.0 79.1 49.58 3.02 0.19 BLIFT 59.2 79.3 50.38 3.05 0.40
 Table 11: Ablation on using comments and/or perception signals from Salicon
-3.2 Discussion
-Tables 1, 2, and 3 contain the results for the visual question answering tasks, Tables 4, 5, 6 contain the results for video and image understanding tasks, Tables 7 contains the results for dense-captioning, and Table 8 contains the results for image and video memorability benchmarks.
-A common trend we observe across all the results is that Behavior-LLaVA performs better than the base model LLaMA-Vid and the finetuned model Ad-LLaVA on all tasks, especially in zero-shot settings.
-Interestingly, the performance gains remain even after fine-tuning on the task dataset (Tables 4, 5, 6).
-The performance gains are smaller for the more low-level tasks of action and object recognition (Tables 1, 2, and 3), but much higher for the more high-level tasks of emotion understanding, sentiment analysis, persuasion strategy classification, and memorability simulation.
-This indicates that receiver behavior has richer signals for higher-level tasks.
-The gains are observed across both image and video benchmarks.
-We also observe that classification using a story generated by GPT-3.5 (following Bhattacharyya et al.
-[13]) results in better performance than only using the video (Tables 1, 2, 4).
-Figures 4, 5, 7, and 3, 6, 8, show several randomly sampled qualitative examples for dense captions generated by Behavior-LLaVA over images and videos respectively.
-It can be noticed that despite not being explicitly trained for this task, the model performs quite well, picking up various artistic, cognitive, and object and material properties.
-From Table 7, while Behavior-LLaVA shows a decrease in correctness over LLaMA-Vid, it shows significant improvement in other aspects, including detail and quality.
-On these aspects, it even comes close to 2.5X larger models (LLaVA-1.6 (34B)).
-Next, we compare the signals from behavioral data of perception and action.
-For this, we compare Behavior-LLaVA trained on the perception data of Salicon using two instruction formats with Behavior-LLaVA trained on BLIFT.
-Further, within BLIFT, we compare the learnings from the task of predicting likes/views with the task of predicting only comments.
-Table 9 contains the results.
-It can be noted from the table that training on just Salicon results in a performance decrease for the lower-level task of action recognition on MSRVTT-QA but improves performance on the higher-level task of Emotion recognition.
-However, the gains are smaller than those observed with training on BLIFT.
-4 Conclusion
-In this paper, we explore the idea of learning behavior leading to learning content better.
-Humans produce behavior in response to content.
-Hence, logically, behavior should contain signals about content, which, if used as a training task, should help in learning content better.
-We follow this line of thought and show that training large vision and language models on user behavior data of comments and likes collected from Reddit and YouTube leads to performance improvements across a wide variety of tasks.
-The gains are higher on higher-level tasks such as emotion recognition, persuasion strategy classification, and question answering and smaller on lower-level tasks like action and object recognition.
-Further, the gains remain even after fine-tuning the VLMs on those benchmarks, thus demonstrating the importance of learning behavior in understanding content better.
-Appendix A GPT-4 Prompts
-Listing 4: GPT-4V Prompt to calculate correctness of a image dense caption
-You are a great critique for analyzing images and captions.
-Assess the performance of a dense image captioning model based on the correctness of the captions generated.
-Please assess the correctness of the provided caption in relation to the image.
-Consider whether the caption accurately identifies and describes the main subjects or objects depicted in the image.
-Assess whether the caption correctly interprets the relationships between elements within the image, such as actions, interactions, or spatial arrangements.
-Focus on the precision and accuracy of the information presented in the caption.
-Provide a score reflecting the level of correctness, ranging from 1 (low correctness) to 10 (high correctness).
-Listing 5: GPT-4V Prompt to calculate detail of a image dense caption
-You are a great critique for analyzing images and captions.
-Assess the performance of a dense image captioning model based on the detail of the captions generated.
-Evaluate the level of detail captured in the provided caption.
-Consider how well the caption describes specific attributes, features, or aspects of the image, including colors, shapes, textures, sizes, and any other relevant details.
-Assess whether the caption provides comprehensive information about the scene depicted in the image, covering both prominent and subtle elements.
-Pay attention to the depth and specificity of the details conveyed in the caption.
-Provide a score indicating the richness of detail, ranging from 1 (low detail) to 10 (high detail).
-Listing 6: GPT-4V Prompt to calculate quality of a image dense caption
-You are a great critique for analyzing images and captions.
-Assess the performance of a dense image captioning model based on the quality of the captions generated.
-Assess whether the caption is concise yet descriptive, providing meaningful and engaging information about the image.
-Evaluate the captions ability to evoke a clear mental image corresponding to the visual content.
-Additionally, consider if the caption is insightful or imaginative in its description.
-Provide a score reflecting the overall quality of the caption, ranging from 1 (low quality) to 10 (high quality).
+Sampling Ratio Epoch Likes/Views $R^{2}$ Comments Perplexity Performance Base-Model 0 -0.1 6.22 0 1:1 0.5 0.11 4.71 5.49 1 0.22 3.95 8.23 1.25 0.33 3.19 10.97 1.5 0.35 3.13 11.79 2 0.38 3.08 12.31 2.2 0.4 3.05 12.57 1:2 0.5 0.14 4.33 3.04 1.05 0.28 3.66 7.08 1.45 0.42 2.99 8.12 1:10 0.5 0.15 3.43 1.38 0.8 0.31 2.78 3.44 1 0.38 2.46 4.48 1.2 0.46 2.13 5.51 2:1 0.5 0.1 5.3 3.52 1 0.21 4.6 7.45 1.5 0.31 3.9 9.13
+Table 12: Ablation on different sampling ratios and epochs of training.
+Sampling ratio is the ratio of behaviour data to multimodal instruct data.
+Performance is the average increase in 0-shot accuracy on 6 tasks with 250 samples each from the eval set.
+These tasks include image emotion recognition, video emotion recognition, persuasion strategy classification, MSRVTT, HVU and MSVD-QA
+Model Audio Summarization (3 Shot) IMDb Sentiment BLEU ROUGE METEOR 0-shot 1-shot Behaviour-LLaVA 19.0 25.1 39.3 84.1 90.2 LLaMA-VID 15.1 18.3 30.7 80.3 87.9 VALOR (Chen et al., 2023) 6.6 10.0 23.9 - - Improvement of Behavior-LLaVA over LLaMA-Vid 25% 37% 28.01% 4.73% 2.61%
+Table 13: Evaluation on audio and text modalities. We evaluate on the audio summarization benchmark(Han et al., 2023) for audio and IMDB sentiment benchmark for text.(Maas et al., 2011)
+Figure 11: Percentage performance improvement over an untrained LLaMA-Vid model, compared across various sampling ratios at different checkpoints.
+The 1:1 sampling ratio shows the best empirical performance.
+Performance is averaged over 0-shot accuracy improvements on six tasks with 250 samples each from the evaluation set.
+These tasks include image emotion recognition, video emotion recognition, and persuasion strategy classification, MSRVTT, HVU and MSVD-QA.
+The figure also shows the benefit of our data filtering process.
+While training on unfiltered BLIFT also improves the result over baseline performance of Llama-Vid but data filtering adds more improvement on top of it.
+Appendix A Dataset Descriptions
+  1. 1.
+    MSVD-QA and MSRVTT-QA: These datasets are based on Microsoft Research Video Description (Chen & Dolan, 2011) and MSR-VTT corpora (Xu et al., 2016b) and are extensively used in many video captioning and question-answering experiments.
+The MSVD-QA dataset has a total number of 1,970 video clips and 50,505 question-answer pairs.
+The MSRVTT-QA dataset contains 10K video clips and 243k question-answer pairs.
+  2. 2.
+    ActivityNet-QA (Caba Heilbron et al., 2015) is a benchmark primarily for human activity understanding containing 849 hours of video, including 28,000 action instances.
+  3. 3.
+    VideoEmotion-8 (Asur & Huberman, 2010) dataset comprises 1,101 user-generated videos sourced from YouTube and Flickr, each containing a minimum of 100 videos per emotional category, as per Plutchik Wheel’s emotion model.
+  4. 4.
+    Ekman-6 (Xu et al., 2016a) dataset is compiled from social websites, with each of its 1,637 videos labeled with a single emotion category based on Ekman’s psychological research.
+  5. 5.
+    CAER (Lee et al., 2019) dataset, sourced from TV shows, consists of 13,201 clips with an average sequence length of 90, each manually labeled with six basic emotions, aligning with the Ekman-6 dataset.
+  6. 6.
+    IAPSa (Mikels et al., 2005) is a subset of IAPS, following the Mikels model with eight emotion categories such as amusement, awe, contentment, excitement, anger, disgust, fear, and sadness.
+It consists of 395 affective images, marking the first visual emotion dataset with discrete categories.
+  7. 7.
+    Emotion6 (Peng et al., 2015) features 1,980 images sourced from Flickr, each labeled by 15 annotators according to the Ekman model, covering six emotion categories: happiness, anger, disgust, fear, sadness, and surprise.
+  8. 8.
+    EmoSet (Yang et al., 2023) encompasses a total of 3.3 million images, including 118,102 from social networks and artistic sources, evenly distributed across various emotion categories.
+Based on the Mikels model, EmoSet is categorized into eight emotion categories.
+  9. 9.
+    Abstract (Machajdik & Hanbury, 2010) exclusively consists of color and texture combinations without recognizable objects.
+Differing from the IAPS dataset where emotions often stem from identifiable objects, the abstract paintings dataset was peer-rated via a web survey, with each image rated approximately 14 times.
+It comprises 228 images spanning eight categories similar to those in IAPS.
+  10. 10.
+    Memento10k(Newman et al., 2020) is a short-term video memorability dataset comprising 10,000 video clips, with 900,000 human memory annotations recorded at various delay intervals.
+The video clips were, on average, 3s long.
+  11. 11.
+    VideoMem (Cohendet et al., 2019) is comprised of 10,000 soundless videos, each lasting 7 seconds, accompanied by memorability scores.
+Memorability was measured twice: first, shortly after viewing and again 24-72 hours later to capture both short-term and long-term memorability effects.
+  12. 12.
+    LaMem (Khosla et al., 2015) dataset is a short-term image memorability dataset comprising of 60000 images.
+The dataset contains scene-centric images, object-centric images and other types such as images of art, images evoking certain emotions, and other user-generated images.
+  13. 13.
+    SUN (Isola et al., 2011) dataset is a short-term image memorability dataset comprising $2$ images that were sourced from the SUN database.
+  14. 14.
+    MemCat (Goetschalckx & Wagemans, 2019) dataset is a short-term image memorability dataset comprising 10000 images.
+It consists of five broader memorability-relevant semantic categories (animal, sports, food, landscapes, vehicles), with 2K exemplars each, further divided into different subcategories (e.g., bear, pigeon, cat, etc.
+for animal).
+The images were sourced from existing image sets: ImageNet, COCO, Open Images Dataset, and SUN.
+  15. 15.
+    MediaEval (Kiziltepe et al., 2021) utilized publicly available links to short-form video clips, each averaging 6 seconds in duration, with both short-term and long-term memorability scores.
+Short-term memorability evaluations were conducted on videos viewed within the preceding few minutes, while long-term memorability assessments were based on videos viewed within the previous 24 to 72 hours.
+  16. 16.
+    LAMBDA (I et al., 2024) is a long-term memorability consisting of 2205 video ads collected over 1749 participants covering 276 brands.
+The average video length is 33 seconds and the videos are highly complex, consisting of audio, logos, fast-moving scenes, emotions, etc.
+Appendix B Limitations
+In this paper, we try to show the hypothesis that training on the behavior modality improves learning of the content modality.
+We train the models on comments and likes to show this.
+We test our models on multiple benchmarks and obtain positive results.
+While we try to cover a wide variety of tasks and while results do conclusively show that the hypothesis is true, yet, we can test on more benchmarks covering even more tasks.
+Similarly, we can show it with multiple models, other than LLaMA-Vid.
+Appendix C Broader Impacts
+Our paper talks about how behavioral training can positively impact content understanding of VLMs.
+We think this will be useful in various content understanding applications such as question answering, captioning, etc.
+Appendix D Ethical Implications
+This paper demonstrates that training on behavioral modalities enhances the learning of content modalities.
+Models trained on user interactions such as comments and likes were tested on multiple benchmarks and yielded positive results.
+While these findings present exciting opportunities for advancing content understanding in AI systems, they also raise important ethical considerations that must be carefully addressed.
+1.
+No personally identifiable information (PII) is used to improve content understanding.
+Instead, aggregated behavioral data, including replays, likes, and comments, is utilized, ensuring user privacy.
+We have implemented rigorous anonymization and aggregation techniques to protect individual user identities.
+2.
+We explicitly acknowledge the inherent biases that may exist in data sourced from social media platforms such as Reddit and YouTube.
+Despite our rigorous data filtering and cleaning processes, we recognize that the dataset may still be subject to demographic skews, self-selection bias, and algorithmic influences.
+These biases could potentially lead to uneven model performance across different user groups or reinforce existing societal biases.
+To mitigate these issues, we emphasize the importance of considering these broader implications when applying our model and interpreting its results.
+3.
+We acknowledge the need for greater cultural diversity in our dataset.
+To address this, we plan to release our artifacts as open-source and encourage community contributions to incorporate multilingual and multicultural data.
+This could involve expanding the range of subreddits and YouTube channels included in our dataset, with the aim of capturing a more diverse and representative spectrum of receiver behavior across different cultural contexts.
+By taking this approach, we hope to enhance the ethical considerations and societal impact of our work, providing a more holistic view of behavioral patterns in various cultural settings.
+However, we also recognize that this approach may introduce new challenges, such as ensuring the quality and reliability of community-contributed data.